### v1

# TripleMixer: A 3D Point Cloud Denoising Model for Adverse Weather

Abstract
LiDAR sensors are crucial for providing high-resolution 3D point cloud data in autonomous driving systems, enabling precise environmental perception. However, real-world adverse weather conditions, such as rain, fog, and snow, introduce significant noise and interference, degrading the reliability of LiDAR data and the performance of downstream tasks like semantic segmentation. Existing datasets often suffer from limited weather diversity and small dataset sizes, which restrict their effectiveness in training robust models. Additionally, current deep learning denoising methods, while effective in certain scenarios, often lack interpretability, complicating the ability to understand and validate their decision-making processes. To overcome these limitations, we introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions: rain, fog, and snow. These datasets retain the original LiDAR acquisition information and provide point-level semantic labels for rain, fog, and snow. Furthermore, we propose a novel point cloud denoising model, TripleMixer, comprising three mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer. These layers are designed to capture geometric spatial information, extract multi-scale frequency information, and enhance the multi-channel feature information of point clouds, respectively. By leveraging the combined capabilities of these three core mixer layers, the TripleMixer network not only effectively preserves and enhances critical features of point cloud data during the denoising process but also offers strong interpretability. Experiments conducted on the WADS dataset in real-world scenarios, as well as on our proposed Weather-KITTI and Weather-NuScenes datasets, demonstrate that our model achieves state-of-the-art denoising performance. Additionally, our experiments show that integrating the denoising model into existing segmentation frameworks enhances the performance of downstream tasks. These results highlight our model’s potential to significantly improve the accuracy and reliability of autonomous driving systems in complex environments. The datasets and code will be made publicly available at https://github.com/Grandzxw/TripleMixer.
Index Terms:
LiDAR Dataset, Weather-KITTI, Weather-Nuscenes, Point Cloud Denosing, Adverse Weather
I Introduction
Figure 1: Visualization of semantic 3D points from the same LiDAR frame across four weather conditions in our proposed Weather-KITTI dataset. Weather noises are marked in red.
LiDAR sensors play a crucial role in autonomous driving systems by emitting laser beams and capturing the reflected signals to generate high-resolution 3D point cloud data, thus providing precise environmental perception[1, 2, 3, 4]. However, the LiDAR data included in current large-scale autonomous driving datasets are typically ”clean”. In real-world scenarios, varying weather conditions, such as rain, fog, and snow, can cause laser beams to scatter, be absorbed, or reflect abnormally, leading to significant noise and interference[5, 6, 7]. This increase in noise not only reduces the reliability of the LiDAR data but also potentially degrades the accuracy of downstream tasks, such as semantic segmentation and object detection, which in turn can affect the overall decision-making capability of autonomous driving systems[8, 9, 10, 11]. Therefore, effectively denoising LiDAR data while preserving its accuracy under diverse weather conditions has become a critical challenge in autonomous driving systems.
To address this challenge, some researchers have specifically developed LIDAR datasets under adverse weather conditions. These datasets can be broadly categorized into two types: real-world datasets, collected under adverse weather conditions, and synthetic datasets, generated through computer simulations[12, 13, 14]. Real-world datasets often face challenges related to the complexity of weather annotation, high data collection costs, and small dataset sizes, which restrict their applicability across a wide range of weather conditions. On the other hand, synthetic datasets allow researchers to create extreme weather scenarios under controlled conditions, enabling more comprehensive testing of system reliability. However, existing synthetic datasets suffer from limitations such as limited weather types, the loss of original LiDAR acquisition information and a lack of precise point-level weather labels.
In addition to these datasets, several denoising methods for point clouds have been proposed, which can be broadly categorized into statistical approaches and deep learning methods[12, 13]. Statistical approaches filter noise by analyzing the distribution patterns of point clouds. However, despite their effectiveness on smaller datasets, they often require substantial computational resources, especially when dealing with large-scale point clouds[15, 16, 17, 18, 19]; In contrast, deep learning methods leverage the powerful feature-learning capabilities of neural networks to directly learn noise patterns from input point clouds. Existing deep learning denoising networks are typically further classified into two categories based on the input format of the point clouds: range image-based methods and 3D point cloud-based methods[20, 21, 22, 23]. The former involves projecting 3D point cloud data onto a 2D range images, but this projection process may lead to a loss of spatial information, potentially affecting the final denoising performance, particularly when dealing with complex point cloud data. The latter directly processes raw 3D point cloud data, better preserving the spatial structure and geometric information, which results in more effective denoising. However, while these deep learning methods have achieved promising results in denoising tasks, they often lack interpretability.
To address the challenges posed by existing datasets and denoising methods, this paper presents two key solutions. First, we introduce two large-scale adverse weather datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions, including rain, fog, and snow. These datasets retain the original LiDAR acquisition information, such as intensity, and provide point-level labels for rain, fog, and snow. Our datasets not only better reflect the performance of LiDAR sensors under different weather conditions in the real world but also provide researchers with more precise and detailed annotations, supporting deeper algorithm research and model optimization. This design aims to provide robust and reliable data support for improving the robustness and reliability of autonomous driving systems under complex weather conditions. Second, we propose a novel point cloud denoising model, the TripleMixer network, which consists of three mixer layers: the Geometry Mixer Layer, Frequency Mixer Layer, and Channel Mixer Layer. The Geometry Mixer Layer focuses on capturing the geometric spatial information of point clouds by analyzing and mixing the geometric relationships of neighboring points, thereby preserving local structural features and important geometric details during the denoising process. The Frequency Mixer Layer is dedicated to extracting frequency information from point clouds by employing multi-scale analysis. This process effectively learns the distinct spectral features from low-frequency components, which represent the global structure, and from high-frequency components, which contain both details and noise, improving the accuracy of point cloud interpretation in complex environments. The Channel Mixer Layer enhances the multi-channel feature information of point clouds by facilitating information exchange and integration between different feature channels, capturing contextual relationships between features, and further improving denoising accuracy. The collaborative effect of these three core mixer layers enables the TripleMixer network to not only preserve and enhance critical features of point cloud data during the denoising process but also provides strong interpretability. Experiments conducted on the WADS dataset[17] in real-world scenarios, as well as on our proposed weather-KITTI and Weather-Nuscene datasets, demonstrate that our model achieves state-of-the-art denoising performance. Additionally, when applying our denoising model to downstream semantic segmentation tasks, experimental results indicate that integrating our denoising model into existing segmentation models can notably improve their segmentation performance. Our main contributions are summarized as follows:
  • We introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, specifically designed for autonomous driving in three adverse weather conditions: rain, fog, and snow. These datasets retain the original LiDAR acquisition information, including intensity, and provide precise point-level weather labels. We believe that these datasets will support efforts to improve the robustness and reliability of autonomous driving systems under complex weather conditions.
  • We propose a novel point cloud denoising model, the TripleMixer network, which consists of three distinct mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer. These layers effectively preserve geometric spatial information, separate frequency components, and enhance multi-channel feature integration. Our TripleMixer network not only preserves and enhances critical features of point cloud data during the denoising process but also provides strong interpretability.
  • Extensive experiments conducted on both existing and newly proposed datasets demonstrate the state-of-the-art denoising performance of the proposed TripleMixer model, and show that integrating our denoising model into existing semantic segmentation frameworks leads to notable improvements in downstream tasks. These findings highlight the practical applicability and effectiveness of the proposed model in enhancing the robustness and reliability of autonomous driving systems in real-world scenarios.
II Related Work
II-A Adverse Weather Datasets
Weather changes are common in the physical world of autonomous driving. Recent datasets investigating LiDAR performance in adverse weather conditions are generally divided into two categories: real-world recordings and synthetic point clouds. Existing real-world datasets face challenges with complex manual weather labeling and limited scale. Meanwhile, current synthetic datasets often fail to retain the original characteristics of LiDAR.
Real-world datasets are typically collected either outdoors or in weather chambers. The DENSE dataset[24] was recorded within fog chambers, which allow for full control over weather conditions and surroundings. However, despite this level of control, these chambers do not adequately reflect real-world conditions. The CADC dataset[25] was recorded during the wintertime in adverse Canadian weather with the VLP32 LiDAR sensors. The dataset has 3D bounding box annotations available, but does not come with weather-related labels and it only covers snowfall weather conditions. The Boreas dataset[26] was collected by driving a repeated route under adverse weather conditions such as rain and falling snow, using a 128-channel LiDAR. It provides 3D bounding box annotations for 3D object detection, but also does not offer weather-related labels. The WADS dataset[17] is an extreme weather dataset from open-world scenarios and includes point-wise weather labels. However, due to the complexity of manually labeling snow noise points individually, it features a limited scale and few scenarios.
Synthetic datasets are collected by augmenting existing datasets. The SnowKITTI dataset[21] is a synthetic dataset based on the KITTI odometry benchmark dataset[27]. However, it only contains data from snowfall weather conditions and has altered the original KITTI point cloud format, with each point cloud frame formatted as 1048×64. Additionally, it has lost the intensity information of each point cloud. The SMART-Degradation dataset[28] provides additional raw data collected by a 32-channel 3D LiDAR and synthesized rainy LiDAR scans. However, it does not provide point-wise weather labels for the rainy LiDAR scans, making it challenging to apply to point cloud perception tasks.
II-B Point Cloud Denoising Method
Adverse weather conditions can impair the effectiveness of LiDAR sensors by introducing noises into their measurements. These noises can significantly degrade the performance of tasks such as perception and mapping[29]. Consequently, denoising point clouds under such conditions is essential to restore measurement clarity. The emerging field of denoising LiDAR point clouds in adverse weather conditions has been explored in only a few studies so far. The primary methodologies employed include statistical approaches and deep learning approaches.
Statistical approaches filter noise by analyzing the distribution patterns of point clouds. Statistical Outlier Removal (SOR) and Radius outlier removal (ROR) [16] filter out noise by analyzing the density of the nearest neighbor points. SOR removes a noisy point by comparing the mean and standard deviation of the distance to its k-nearest neighbors. ROR processes each point in the point cloud, counting the number of neighbors within a specified search radius to remove noise points. Dynamic ROR (DROR) [30] enhances the ROR filter by dynamically adjusting the search radius. Dynamic SOR (DSOR) [17] is an extension of SOR filter designed to address the non-uniform density within point clouds. It dynamically calculates a local distance threshold to remove noise points. Low-Intensity Outlier Removal (LIOR)[19] and Dynamic Distance–Intensity Outlier Removal (DDIOR)[18] integrates the distance and intensity of points based on the systematic and accurate analysis of data characteristics in snowy weather. Conditional Random Fields Outlier Removal (CRFOR)[31] proposes a de-snowing formulation with Conditional Random Fields and Spatio-Temporal information. Since statistical approaches typically remove noise through iterative algorithms, they are computationally expensive.
Contrary to statistical approaches, deep learning approaches became popular due to the model’s abilities to directly understand the characteristics of weather-induced noise in point cloud. Some general-purpose point cloud semantic segmentation approaches [32, 33, 34] have been adapted to denoise point clouds affected by adverse weather. Although they have achieved good performance, they are not specifically designed for denoising. Existing denoising networks are typically categorized based on the input format of the point cloud: range images and raw 3D point clouds. WeatherNet[20] uses modified LiLaBlock[35] to filter out adverse weather effects in point cloud data. 4DenoiseNet[21] proposes a k-nearest neighbor convolution kernel that captures spatial-temporal information, and with a attention mechanism. However, its limitation lies in the requirement for LiDAR scans with temporal continuity. LiSnowNet[36] combines the Fast Fourier Transform (FFT) with Convolutional Neural Networks (CNNs) for unsupervised de-noising to remove snow from LiDAR point clouds. The denoising networks mentioned above all use range projection images as inputs. While range projection methods usually benefit from fast computations, they face limitations such as projection errors and loss of spatial information when converting 3D point clouds into 2D range images. 3D-OutDet[22] takes raw 3D point clouds as inputs and develops a NeighborHood (NH) Convolution for unorganized scattered 3D point clouds, specifically targeting the removal of snowflakes.
Table I: Comparison of the Proposed Adverse Weather Dataset with Existing Public Datasets. Including: Year, LiDAR Parameters, Number of Dataset Frames, Presence of Semantic Labels, Types of adverse Weather Covered,
Inclusion of LiDAR Intensity Features, and Presence of Point-Wise Weather Labels.
Dataset Year LiDAR Size Semantic Labels Weather Types Intensity Weather Labels
Snow Fog Rain
WADS[17] 2021 64 channel 1K frames point-wise
CADC[25] 2021 32 channel 7K frames bounding boxes
DENSE[24] 2020 32 & 64 channel 13K frames bounding boxes
Boreas[26] 2023 128 channel 7K frames bounding boxes
SMART-Degradation[28] 2021 32 channel 1K frames -
Oxford RobotCar[37] 2020 32 channel 11K frames -
SemanticKITTI[38] 2019 64 channel 2K frames point-wise
nuScenes-lidarseg[39] 2020 32 channel 28K frames point-wise
SnowKITTI[21] 2023 64 channel 43K frames -
Weather-KITTI (Ours) 2024 64 channel 130K frames point-wise
Weather-NuScenes (Ours) 2024 32 channel 84K frames point-wise
III Datasets Generation
As introduced in the previous section, current adverse weather datasets face several challenges: Firstly, there is an absence of point-wise semantic labels and weather labels; Secondly, manually annotating weather conditions in large-scale, real-world datasets is highly complex; Thirdly, existing synthetic datasets lack the original LiDAR acquisition information such as intensity, which impacts the performance of LiDAR downstream perception tasks. Building on this, we propose our synthetic adverse weather datasets, named Weather-KITTI and Weather-NuScenes, which are based on the SemanticKITTI[38] and nuScenes-lidarseg[39] datasets, respectively. To simulate three adverse weather conditions: snow, fog, and rain, we employ the LiDAR Snowfall Simulation (LSS) [40] for snow, the LiDAR Fog Simulation (LFS) [41] for fog, and the LiDAR Light Scattering Augmentation (LISA) [42] for rainy conditions. Table I presents the characteristics of the proposed dataset and existing publicly datasets. Compared to these existing adverse weather datasets, our proposed dataset provides a larger scale, covers a broader range of weather scenarios, offers complete original point cloud information, and has the capability to provide point-wise semantic and weather labels.
III-A Adverse Weather Simulator
LiDAR sensors emit laser pulses and measure travel distance based on the echoes received from targets, thereby capturing the precise shape of the targets[43]. However, the LiDAR pulse system is susceptible to interference from scattering media such as particles of rain, fog, and snow, which can lead to shifts in measured distances, variations in echo intensity, and missing points, among other issues[44, 45, 46].
The received signal power at a LiDAR’s receiver can typically be described using a linear system model[43], which is valid for non-elastic scattering. In this model, the range-dependent received signal power $P_{R}$ is modeled as the time-wise convolution between the time-dependent transmitted signal power $P_{T}$ and the time-dependent impulse response $H$ of the environment:
[…]
where $C_{A}$ is a system constant independent of time $t$, $R_{0}$ is the object distance and $c$ is the speed of light.
1) Snow Model: In this work, we follow [40] to simulate these snow-corrupted data. This physically-based method samples snow particles in the 2D space and modify the measurement for each LiDAR beam in accordance with the induced geometry. For each point $p(x,y,z,i)$ from LiDARs, we calculate its snow-attenuated point $p_{\text{snow}}$ as follows:
$R^{*}=\arg\max\left({P}_{R}\right)-c\frac{\tau_{H}}{2},\$ (2)
$i_{\text{snow}}=P_{R}+i_{\max}f_{s}\left|f_{o}-\left(1-\frac{R^{*}}{R_{\max}}\right)\right|^{2},\$ (3)
${{p_{\text{snow}}}}=\left(\frac{R^{*}}{R_{0}}\cdot(x,y,z),i_{\text{snow}}\right),$ (4)
where $\tau_{H}$ is the LiDAR sensor’s half-power pulse width. $R^{*}$ and ${R_{\max}}$ represent the distances to the object at maximum received power $P_{R}$ and the furthest detectable object, respectively. $f_{s}$ is the focal slope and $f_{o}$ is the focal offset of the LiDAR sensor. As the snowfall rate $r_{s}$ increases, so does the number of particles in a point cloud frame, leading to denser snow particle accumulations in heavier snowfalls. Similar to [17], we define point clouds with changed spatial positions as the snow class.
2) Fog Model: The physically-based fog simulation[41] is employed to transform clear LiDAR point clouds into their foggy counterparts. For each original point cloud, we calculate its fog-attenuated point $p_{\text{fog}}$ as follows:
$i_{\text{hard }}={i}\cdot\exp({-2\alpha R_{0}}),$ (5)
$i_{\text{soft}}={i}\cdot\frac{R_{0}^{2}}{\beta_{0}}\beta\cdot{i}_{tmp},$ (6)
${{p_{\text{fog}}}}=\left\{\begin{array}[]{ll}\left(s\cdot n\cdot(x,y,z),{i_{\text{soft }}}\right),&\text{ if }{i_{\text{soft }}}>{i_{\text{hard }}}\\ \left(x,y,z,{i_{\text{hard }}}\right),&\text{ else. }\end{array}\right.$ (7)
where $\alpha$ is the attenuation coefficient, $\beta$ denotes the fog backscattering coefficient, $\beta_{0}$ describes the differential reflectivity of the target. $i_{\text{soft }}$ and $i_{\text{hard }}$ represent the received responses for the soft target term and the hard target term, respectively. $s$ denotes scaling factor and $n$ is noise factor. ${i}_{tmp}$ is the maximum response of the soft-target term. As $\beta$ increase, the fog rate rises, and the number of fog particles in the point cloud correspondingly increases. We similarly map point clouds with changes in three-dimensional spatial positions as the fog class.
2) Rain Model: We employ the physics-based simulation model[42] to augment LiDAR scans for rain condition, which is based on a hybrid Monte-Carlo based approach. Specifically, the rain-distorted point cloud ${{p_{\text{rain}}}}$ is calculated as follows:
[…]
${i_{\text{rain}}}={i}\cdot\exp^{-2\alpha R_{\text{max }}}\cdot\min\left\{\left(\frac{D_{\text{max }}}{D_{b}\left(R_{\text{max }}\right)}\right)^{2},1\right\},$ (9)
${{p_{\text{rain}}}}=\left(\gamma\cdot(x,y,z),i_{\text{rain}}\right),$ (10)
where $\alpha$ is the extinction coefficient, $D$ is the particle diameter and $q_{e}$ is the extinction efficiency. $N$ denotes the particle density function. $D_{b}$ represents the beam diameter at the object distance $R_{\text{max }}$ and $\gamma$ is the rain-attenuated coefficient. Similarly, as the rain rate $r_{r}$ increases, the number of rain particles in a point cloud frame also correspondingly increases. We map point clouds with a rain-attenuated coefficient $\gamma$ not equal to 1 as the rain class.
((a))
((b))
Figure 2: LiDAR data frames for different levels of weather severity in our proposed Weather-KITTI and Weather-NuScenes datasets. LiDAR frames corresponding to different weather severities are randomly selected.
((a))
((b))
Figure 3: Semantic distribution of 3D points for all sequences in Weather-KITTI and Weather-NuScenes.
Table II: Details of Weather Simulations. Each Weather
Simulation Is Categorized Into Light, Moderate,
and Heavy Levels Based on Different
Weather Parameters.
Weather Intensity Description
Snowfall Simulation Light snowfall rate $r_{s}=[0.5,1.0]$ mm/h
Moderate snowfall rate $r_{s}=[1.5,2.0]$ mm/h
Heavy snowfall rate $r_{s}=[2.5,3.0]$ mm/h
Fog Simulation Light fog simulation with $\beta=[0.01,0.05]$
Moderate fog simulation with $\beta=[0.08,0.14]$
Heavy fog simulation with $\beta=[0.18,0.25]$
Rain Simulation Light rain rate $r_{r}=[1.0,1.5]$ mm/h
Moderate rain rate $r_{r}=[1.8,2.4]$ mm/h
Heavy rain rate $r_{r}=[2.6,3.0]$ mm/h
III-B Dataset Statistics
Adverse weather primarily causes back-scattering and attenuation of LiDAR pulse transmissions, leading to significant shifts in both range and intensity for points in a LiDAR point cloud[43, 47]. Figure 1 illustrates the semantic distributions of the same point cloud frame under three different weather conditions within our proposed Weather-KITTI dataset and their corresponding clean datasets. The weather noise points induced by the three weather conditions are primarily concentrated around the LiDAR sensor. Snowflakes contain ice crystals with varying densities, and in snowy conditions, each laser beam interacts with airborne particles, reflecting at different angles and considering potential occlusions. Compared to rain and fog, which consist of denser small water droplets, snowflakes result in more dispersed noise points.
Table II delineates the severity levels of various weather-induced corruptions as configured in this study. In the snow simulation, the snowfall rate $r_{s}$ is set at [0.5, 1.0], [1.5, 2.0], and [2.5, 3.0] mm/h to represent light, moderate, and heavy snowfall, respectively. For fog simulations, the fog parameter $\beta$ is adjusted to [0.01, 0.05], [0.08, 0.14], and [0.18, 0.25] to reflect light, moderate, and heavy fog conditions. Similarly, in the rain simulation, the rain rate $r_{r}$ is set to [1.0, 1.5], [1.8, 2.4], and [2.6, 3.0] mm/h to simulate light, moderate, and heavy rainfall. As the severity of the weather increases, the extent of point cloud corruption proportionally intensifies. According to Table II, all LiDAR frames of Weather-KITTI and Weather-NuScenes are randomly synthesized with three different levels of weather severity in this study. This configuration allows us to investigate the generalizability of during models training. Simultaneously, we ensure that each level of weather severity has a similar number of LiDAR frames, guaranteeing the diversity of data across different weather scenarios. As depicted in Figure 2(a), the number of LiDAR frames for each weather category under the three weather severity levels in Weather-KITTI is approximately 14K frames. Similarly, as shown in Figure 2(b), the number of LiDAR frames for each weather category under each weather severity level in Weather-NuScenes is around 9K frames.
Figure 4: Overview of the proposed denoising model, the TripleMixer network. The TripleMixer network comprises three main layers: the Geometry Mixer (GMX) Layer, the Frequency Mixer (FMX) Layer, and the Channel Mixer (CMX) Layer. These layers are detailed in Section IV.
Weather-KITTI is a large-scale synthetic LiDAR dataset designed for adverse weather conditions, based on SemanticKITTI [38]. It includes 22 sequences with a total of 130,656 LiDAR scans. Each point cloud is segmented into 21 semantic classes, with the semantic classes for snow, fog, and rain mapped to labels 110, 111, and 112, respectively, across all sequences. As SemanticKITTI provides semantic labels only for the first ten sequences, we also limit the availability of other semantic labels to the first ten sequences. Weather-NuScenes, derived from nuScenes-lidarseg [39], contains 84,390 LiDAR scans evenly divided into ten sequences. It includes 19 semantic classes and has similarly been enhanced with semantic classes for snow, fog, and rain across all sequences, mapped to labels 110, 111, and 112, respectively. Figure 3 illustrates the distribution of LiDAR points across semantic categories for all sequences in both Weather-KITTI and Weather-NuScenes. As depicted in the Figure 3(a), Weather-KITTI contains 139 million, 383 million, and 89 million LiDAR points for snow, fog, and rain noise, respectively. In comparison, as shown in Figure 3(b), Weather-NuScenes has 23 million, 112 million, and 6 million LiDAR points for snow, fog, and rain noise, respectively. This highlights the large scale and diverse weather conditions captured in our proposed datasets.
IV Denoising Network
IV-A Problem Statement
In this study, we aim to perform adverse weather denoising from LiDAR point clouds to restore clean point clouds. Given an adverse weather dataset of $N$ LiDAR scans $\left\{\left(P_{i},L_{i}\right)\mid i=1,\ldots,N\right\}$, where $P_{i}\in\mathbb{R}^{n_{i}\times 5}$ is the $i$th point set containing $n_{i}$ LiDAR points. Each row of $P_{i}$ consists of five features representing one LiDAR point $p$, namely ($x$, $y$, $z$, $i$, $r$). ($x$, $y$, $z$) denote the Cartesian coordinates of the point relative to the LiDAR, $i$ indicates the intensity of the returning laser beam and $r$ represents the 3D spatial distance from the point to the LiDAR center. $L_{i}\in\mathbb{Z}^{n_{i}}$ contains the ground truth labels for each point $p$ in $P_{i}$. Our objective is to learn a denoising model $\Phi$ that outputs a label set $\hat{L_{i}}\in\mathbb{Z}^{n_{i}}$, designed to remove outliers from the point cloud data as follows:
$\hat{L_{i}}=\Phi({P_{i}};{\theta}),$ (11)
where $\theta$ denotes the model parameters to be optimized, aiming to minimize the difference between the prediction $\Phi({P_{i}};{\theta})$ and the ground truth labels $L_{i}$.
IV-B The Overview of Network
In this section, we propose the TripleMixer, a robust denoising model designed for removing 3D point cloud noises under adverse weather conditions. The architecture of our model is depicted in Figure 4. The TripleMixer consists of three primary layers: 1) the Geometry Mixer (GMX) Layer, which mixes the spatial geometric relationships of neighboring points, thereby enabling the network to discern crucial local structural features and maintain important geometric details; 2) the Frequency Mixer (FMX) Layer, employing multi-scale wavelet analysis to achieve a more comprehensive representation of both noise and normal points in the frequency domain, thereby enhances the model’s ability to discern fine-grained characteristics; and 3) the Channel Mixer (CMX) Layer, which enhances the overall feature representation by mixing contextual information across channels, improving the model’s accuracy and adaptability in complex environments. Specifically, in the GMX layer, we first voxelize the LiDAR point cloud in 3D space to facilitate downsampling, then select K neighboring points and employ the attentive pooling operation to effectively capture local geometric patterns. For the FMX layer, we initially quantize the 3D features derived from the GMX layer along the X, Y, and Z axes into 2D grids. These grids are projected onto the YZ, XZ, and XY planes, creating the corresponding 2D grid images for triple plane projection. Subsequently, these images are processed using an adaptive wavelet module that provides a multi-resolution decomposition, efficiently enhancing the learning of multi-scale features. In the CMX layer, we mix feature information across different channels, synthesizing information across various directions and scales, thereby enriching the network with superior contextual information. Detailed discussions of these components will follow in subsequent sections.
Figure 5: Details of the Geometry Mixer Layer. The input features within the same voxel are meaned (downsampled), followed by the KNN neighborhood feature encoding, an attentive pooling operation, and a residual connection to obtain the output features.
IV-C Geometry Mixer (GMX) Layer
The GMX layer is a pivotal component of our TripleMixer, which efficiently encodes and aggregates local geometric features, enhancing the model’s ability to analyze 3D point cloud. By mixing spatial information at the intra-voxel level and geometric information at the inter-point level, the GMX layer captures the spatial structures and relationships around each point, thereby forming a richer and more discriminative feature representation. The structure of the GMX layer is shown in Figure 5. The GMX layer includes Voxel Mixing and Point Mixing. Initially, Voxel Mixing is performed by averaging (downsampling) all point features within the same voxel, followed by an Multi-Layer Perceptron (MLP) to introduce non-linearity, resulting in a unique center point $p_{i}\in\mathbb{R}^{5}$ that best represents the voxel. Then, for each of the nearest K points $\left\{p_{i}^{1},\cdots p_{i}^{k},\cdots p_{i}^{K}\right\}\in\mathbb{R}^{5}$ of a center point $p_{i}$, we explicitly perform local point geometric feature mixing $l_{i}^{k}$ as follows:
$l_{i}^{k}=MLP\left(p_{i}\oplus p_{i}^{k}\oplus\left(p_{i}-p_{i}^{k}\right)\right),$ (12)
where $\oplus$ is the concatenation operation, and $p_{i}-p_{i}^{k}$ represents the geometric feature similarity between the center point $p_{i}$ and the k-th neighboring point $p_{i}^{k}$, the MLP effectively models the spatial relationships between these points by learning from the concatenated features and their geometric similarities. Subsequently, we apply attentive pooling operation to automatically learn and emphasize important local features:
$c_{i}^{k}=Softmax\left(\left(FC(l_{i}^{k})\right)\right),$ (13)
$f_{i}=\sum_{i=1}^{K}c_{i}^{k}\odot l_{i}^{k},$ (14)
where $c_{i}^{k}$ is the unique attention score for the feature $l_{i}^{k}$, and $\odot$ denotes the dot product. In Eq. (13) and Eq. (14), a fully connected layer (FC) is used for point feature encoding, and the Softmax function is employed to learn the score $c_{i}^{k}$. These scores are then used to weight the features, which are subsequently summed to generate a new informative feature $f_{i}$. Finally, to prevent overfitting and preserve foundational geometric features, a residual connection is utilized to integrate the original point geometric features $p_{i}$ with the important local features $f_{i}$ obtained from the attentive pooling. The final output features of the GMX layer is computed as:
$F_{i}=Max\left(\left(MLP(f_{i})\right)\right)\oplus p_{i}.$ (15)
Figure 6: Architecture of the Frequency Mixer Layer. The Frequency Mixer Layer consists of three modules: the Triple Plane Projection (TPP) Block, the Lifting Wavelet (LWL) Block, and the Multi-resolution Feature Mixing (MFM) Block. It accepts input features with dimensions [H, V, C], where H and V represent height and width, and C represents the number of channels.
IV-D Frequency Mixer (FMX) Layer
In this section, we introduce the Frequency Mixer (FMX) Layer that designs a denoising lifting wavelet module to capture multi-scale geometric representation in the frequency domain, with its architecture is illustrated in Figure 6. Noise in point clouds is usually randomly distributed in the spatial domain, manifesting as high-frequency components in the frequency domain, while the main structure of the original point cloud is represented by low-frequency components in the frequency domain. Our proposed denoising lifting wavelet module can decompose the features into different scale sub-bands, each containing various frequency information. By integrating these sub-bands, our model can effectively utilize both high-frequency and low-frequency components, resulting in a comprehensive and robust feature representation.
The FMX layer consists of three blocks: 1) the triple plane projection block, which projects the 3D point cloud onto 2D images. This block simplifies the inherent complexity of processing 3D point clouds while ensuring that sufficient spatial information is retained. 2) the lifting wavelet block, which performs a hierarchical decomposition of each 2D images into four sub-band components. This block efficiently generates both approximation and detail coefficients, capturing both low-frequency and high-frequency information at each scale. 3) the multi-resolution feature mixing block, which mixes decomposed components from multiple scales and combine this information across multiple resolutions to significantly enhance feature representation and analysis.
Triple Plane Projection (TPP) Block . It’s aims to obtain the 2D features $x_{i}$ of the triple projection planes from the 3D features $F_{i}$ outputted by the GMX layer. To address the intensive computational demands of 3D convolution operations and the loss of height-related details in BEV projection, we introduce a novel approach by projecting the 3D point cloud along the X, Y, and Z axes onto three distinct 2D planes at varying resolutions, as illustrated in Figure 6. The TPP block utilizes sparse tensor operations to efficiently handle data. It starts by identifying non-zero indices in the point cloud to create sparse projection matrices for quantizing 3D points into 2D planes. The value of each pixel in the 2D plane represents the average of the 3D features falling within the same pixel. This triple plane projection method reduces redundancy and computational load while capturing detailed object features and enhancing spatial information, leading to more precise and accurate object characteristics.
Lifting Wavelet (LWL) Block. It consists of a horizontal wavelet decomposition (HWD) followed by two independent vertical wavelet decomposition (VWD) that generate the four sub-bands of the input 2D feature from the TPP block at each scale, as illustrated in Figure 6. Suppose that the input 2D feature at decomposition layer $t$ is $x^{t}_{i}\in\mathbb{R}^{H\times V\times C}$, the LWL block first splits the input feature $x^{t}_{i}$ into two non-overlapping partitions along the horizontal direction, divided into even and odd components denoted as $x^{t}_{e}\in\mathbb{R}^{H\times V/2\times C}$ and $x^{t}_{o}\in\mathbb{R}^{H\times V/2\times C}$, respectively. The horizontal wavelet decomposition yields the detail components $d^{t}$ and the approximation components $c^{t}$, the details components $d^{t}$ represent the difference between $x^{t}_{o}$ and a predictor operator $P\left(\cdot\right)$ on $x^{t}_{e}$. Conversely, the approximation components $c^{t}$ are computed as $x^{t}_{e}$ plus an update operator $U\left(\cdot\right)$ on $d^{t}$. The process of the horizontal wavelet decomposition is described as follows:
$d^{t}=x^{t}_{o}-P\left(x^{t}_{e}\right),$ (16)
$c^{t}=x^{t}_{e}+U\left(d^{t}\right),$ (17)
where the structure of both the $P(\cdot)$ and $U(\cdot)$ operators is shown in Figure 6. The operation for each begins with reflection padding to keep the original sequence length. This is followed by a $1\times 3$ convolutional layer to process the features, activated by ReLU for non-linearity. Subsequently, a $1\times 1$ convolution then adjusts the channels, and a Tanh activation function retain negative values in the output.
Next, two independent VWD module, which have a structure similar to the HWD module, are respectively applied to the approximation components $c^{t}$ and the details components $d^{t}$ obtained from the HWD. Consequently, the four sub-band components $\left[{x^{t}_{HL},x^{t}_{HH},x^{t}_{LL},x^{t}_{LH}}\right]\in\mathbb{R}^{H/2\times V/2\times C}$ at layer $t$ can be obtained as follows:
[…]
where $L$ and $H$ denote low and high frequency information, respectively, while the first and second positions in each sub-band component’ indices correspond to the horizontal and vertical directions, respectively. Finally, we concatenate the four sub-bands along the channel dimension, obtaining the output feature $x^{t}_{w}$ of the LWL block:
$x^{t}_{w}=x^{t}_{LL}\oplus x^{t}_{LH}\oplus x^{t}_{HL}\oplus x^{t}_{HH},$ (20)
where $x^{t}_{w}\in\mathbb{R}^{H/2\times V/2\times 4C}$. Unlike pooling or strided convolutions, the LWL block is information lossless, as it expands the number of channels by the same factor by which it reduces spatial resolution. The low frequency sub-band $x^{t}_{LL}$ serves as the input feature for the lifting wavelet decomposition in the next layer $t+1$. This recursive process is repeated at each scale, capturing multi-scales geometric information effectively. The maximum number of scales can be determined analytically based on the initial resolution of the triple projection planes, as detailed in SectionVI-A.
Multi-resolution Feature Mixing (MFM) Block. It comprises an MLP layer, a transposed convolution (TConv), a batch normalization (BN) layer, and residual connections. The concatenated feature $x^{t}_{w}$ from the LWL block is sequentially processed through these modules, as detailed below:
$W^{t}_{i}=BN\left(TConv\left(MLP(x^{t}_{w})\right)\right)\oplus x^{t}_{i},$ (21)
where $x^{t}_{w}$ is initially processed through an MLP to restore the original channel dimension and mix multi-resolution features in the frequency domain. Subsequently, transposed convolutions are employed to upsample the feature maps, restoring the spatial resolution of the sub-band feature maps and providing access to both local and global information. The batch normalization (BN) layer normalizes the feature maps, stabilizing the training process and enhancing training stability and generalization. Finally, a residual connection reintroduces the original input feature $x^{t}$ at layer $t$, ensuring that while complex features are extracted, essential information from the initial input is retained. This process is repeated at multiple scales within the FMX layer, recursively mixing and enhancing multi-scale geometric representations.
IV-E Channel Mixer (CMX) Layer
The CMX layer plays a crucial role in reprojecting the 2D features processed by the FMX layer back into the 3D space, mixing contextual information across channels, and enhancing the overall feature representation. The structure of the CMX layer is shown in Figure 4. This layer is essential for integrating multi-scale geometric features $W_{i}$ from the FMX layer are effectively combined and utilized for subsequent point denoising tasks. The processing flow of the CMX layer is as follows:
$C_{i}=Drop\left(GConv\left(MLP\left(BN(W_{i})\right)\right)\right)\oplus x_{i},$ (22)
where $Drop$ represents the dropout operation to prevent overfitting. $GConv$ represents the group convolution used to mix channel feature information. Initially, the input feature $W^{t}$ undergoes the BN operation. Then, an MLP layer enhances dimensionality, followed by a group convolutional layer for mixing contextual information. A dropout operation prevents overfitting. Finally, a residual connection reintroduces the original input feature $x_{i}$ across the channels, capturing and fusing local-to-global context.
IV-F Loss Function
To train our proposed model and ensure that the captured information is relevant to the point cloud denoising task, the overall loss function consists of three components: the cross-entropy loss, the lovász loss[48] and the wavelet regularization terms. These components are structured as:
$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\text{lovász }}+\mathcal{L}_{\mathrm{wr}}$ (23)
$\mathcal{L}_{\mathrm{wr}}=\lambda_{1}\sum_{t=1}^{N}\left\|\mathcal{D}_{t}\right\|_{2}^{2}+\lambda_{2}\sum_{t=1}^{N}\left\|\mathcal{A}_{t}-\mathcal{A}_{t-1}\right\|_{2}^{2},$ (24)
where $\mathcal{L}_{\mathrm{ce}}$ represents the cross-entropy loss, assessing the accuracy of noise predictions, and $\mathcal{L}_{\text{lovász}}$ denotes the lovász loss, which is designed for improving performance in tasks with imbalanced classes. $\mathcal{L}_{\mathrm{wr}}$ represents the wavelet regularization terms, which include two term. The first wavelet regularization term minimizes the sum of the $l_{2}$ norm of the detail components across all levels, enhancing detail preservation. The second wavelet term minimizes the $l_{2}$ norm of the difference between the approximation components of consecutive levels across all levels, crucial for preserving the mean of the input signal and ensuring a proper wavelet decomposition enforcing a consistent wavelet decomposition structure, which is essential for effectively capturing multi-scale geometric information. $\lambda_{1}$ and $\lambda_{2}$ are the wavelet regularization parameters used to adjust the strength of the regularization terms. $\mathcal{D}_{t}$ represents the mean of the detail sub-bands at decomposition level $t$, while $\mathcal{A}_{t}$ and $\mathcal{A}_{t-1}$ denote the mean of the approximation sub-bands at levels $t$ and level ${t-1}$ of the lifting wavelet, respectively.
V Experiments
In this section, we first introduce the experimental datasets and their division into training, validation, and testing subsets for model training. Next, we describe the implementation details and evaluation metrics of the experiments. Moreover, we conduct extensive performance comparison experiments to benchmark our model against state-of-the-art models. Finally, comprehensive ablation studies are performed to validate the contribution of each component of our model.
V-A Datasets
We conduct qualitative tests on three large-scale LiDAR datasets: WADS[17], as well as the Weather-KITTI and Weather-NuScenes datasets introduced in this paper. These datasets contain a variety of adverse weather conditions. The details of these datasets will be introduced as follows.
WADS[17] contains LiDAR scans from 20 different sequences, with around 1300 LiDAR frames. The majority of these LiDAR scans are collected in snowy urban driving scenarios. Points in a LiDAR scan has been labeled into one of 22 semantic classes. The default labels for snow points are mapped as 110. We divided the WADS dataset [^1] The training set is: [14,15,18,20,24,28,34,36,37]. The val set is: [11,16]. The test set is: [12,13,17,22,23,26,30,35,76]. into training, validation, and testing sets according to the protocol outlined in[22]. To ensure a fair comparison for the point cloud denoising task under adverse weather conditions, we re-trained state-of-the-art models using our splits.
Table III: Splitting of Training, Validation, and Testing Sets for Our Proposed Weather-KITTI and
Weather-NuScenes Datasets.
Weather-KITTI Weather-NuScenes
Snow Scenarios Train: [00,02,08,17,19] [01,02,07,08]
Val: [04,11,12,16] [04]
Test: [01,03,05,06,07,
09,10,13,14,15,
18,20,21] [00,03,05,06,09]
Fog Scenarios Train: [02,03,04,16,17,
18,19,20] [01,02,03,07]
Val: [06,07,09,11] [04]
Test: [00,01,05,08,10,
12,13,14,15,21] [00,05,06,08,09]
Rain Scenarios Train: [01,02,03,04,06,
12,16,17,21] [04,05,07,09]
Val: [09,19,19] [08]
Test: [00,05,07,08,10,
13,14,15,18,20] [00,01,02,03,06]
Weather-KITTI and Weather-NuScenes contain three different weather scenarios: snow, fog, and rain. Specifically, the Weather-KITTI dataset consists of 22 sequences for each weather scenario, while the Weather-NuScenes dataset includes 10 sequences per scenario. For model training, we divided the training, validation, and testing sets as detailed in Table V-A.
V-B Implement Details
In our experiments, we downsampled the point clouds by retaining only one point per 10 cm voxel. The resolutions for the 2D projections along the X, Y, and Z axes are set to [256, 256, 32], respectively. We optimized the model parameters using the AdamW optimizer[49] for 20 epochs with a weight decay of 0.005. The learning rate scheduler first performed a linear warmup over the initial 2 epochs, gradually increasing the learning rate from 0 to 0.001. After the warmup phase, it transitioned into a cosine annealing phase, where the learning rate decayed following a cosine curve, eventually reaching a minimum of $10^{-5}$ by the end of the last epoch. We set the batch size to 4 for each GPU and employed various data augmentation techniques to enhance model robustness, including random rotations, scaling, and random flipping along the XY plane. Additionally, we used stochastic depth with a layer drop probability of 0.2. The loss is the sum of the cross-entropy loss, the lovász loss and the wavelet regularization terms. The regularization parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 for all the experiments. All results are produced on an Ubuntu 18.04 system equipped with two Intel® Xeon® Platinum 8280 CPUs, two Nvidia V100 GPUs, and 32 GB of RAM.
V-C Evaluation Metrics
Following prior work[22, 21, 20], we evaluate the point cloud denoising task using $precision=\frac{TP}{{TP+FP}}$, $recall=\frac{TP}{{TP+FN}}$, F1 score, and mean Intersection over Union (mIoU) as quantitative metrics. Specifically, the F1 score and mIoU are defined as follows:
${F_{1}}=2\times$ (25)
${mIoU}=\frac{TP}{{TP+FP+FN}}$ (26)
Where $TP$ refers to the correctly identified noise points, $FP$ refers to the non-noise points incorrectly identified as noise, and $FN$ refers to the noise points that were not correctly identified.
V-D Quantitative Result
Results On WADS: We compare our model with the current state-of-the-art models, including both statistical and deep learning approaches. Specifically, SOR[16], ROR[16], DSOR[17], and DROR[30] are statistical methods. SalsaNext[32] and Cylinder3D[33] are general point cloud segmentation methods that we have adapted for binary classification to enable comparison. WeatherNet[20], 4DenoiseNet[21], and 3D-OutDet[22] are specialized point cloud denoising networks.
Table LABEL:wads presents a comprehensive comparison of denoising performance on the WADS[17] dataset, evaluating precision, recall, F1 score, and mIoU. The WADS[17] dataset represents real-world conditions, making it a robust benchmark for performance assessment. As illustrated in Table LABEL:wads, while Cylinder3D[33] achieves the highest precision with a score of 97.12, our model closely follows with a minimal difference of approximately 0.06. More importantly, our model outperforms all existing methods in other critical metrics, achieving the highest recall, F1 score, and mIoU with scores of 93.93, 94.97, and 90.73, respectively. These results underscore the superior efficacy of our model in point cloud denoising, particularly in complex real-world applications.
Table IV: Denoising Performance Results on WADS Datasets. The Proposed TripleMixer Achieves the Best mIoU.
Method Precision↑ Recall↑ F1↑ mIoU↑
ROR[16] 17.13 91.80 29.15 17.06
DSOR[17] 65.92 90.93 76.43 61.86
DROR[30] 69.84 90.10 78.68 64.05
SalsaNext[32] 74.16 93.50 82.71 70.52
Cylinder3D[33] 97.12 92.61 94.81 90.13
WeatherNet[20] 96.69 81.24 88.28 79.02
4DenoiseNet[21] 96.46 86.01 90.94 83.38
3D-OutDet[22] 97.10 92.25 94.61 89.78
TripleMixer (Ours) 96.38 93.93 95.13 90.73
Table V: Denoising Performance Results On The Proposed Weather-KITTI Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
Method Average Snow Scenarios Fog Scenarios Rain Scenarios
mIoU Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑
DROR[30] 56.24 44.18 97.36 60.77 43.66 66.05 92.74 78.63 64.79 61.47 96.68 74.11 60.28
SalsaNext[32] 89.58 97.40 92.83 95.06 90.58 97.19 94.59 95.87 92.07 97.75 87.82 92.52 86.08
Cylinder3D[33] 94.28 99.13 96.39 97.74 95.58 99.78 92.92 96.23 92.73 96.79 97.58 97.18 94.52
WeatherNet[20] 79.15 83.28 97.20 89.70 81.33 89.16 86.64 87.88 78.38 78.49 98.81 87.48 77.75
4DenoiseNet[21] 82.85 85.23 97.08 90.77 83.10 89.90 93.95 91.88 84.98 83.02 96.34 89.18 80.48
3D-OutDet[22] 90.55 98.14 95.15 96.63 93.48 99.61 85.26 91.88 84.97 94.12 98.98 96.48 93.21
TripleMixer (Ours) 96.31 99.23 97.18 98.16 96.46 99.03 96.50 97.74 95.61 97.66 99.15 98.40 96.87
Table VI: Denoising Performance Results On The Proposed Weather-NuScenes Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
Method Average Snow Scenarios Fog Scenarios Rain Scenarios
mIoU Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑
DROR[30] 42.93 38.36 92.74 54.27 37.24 40.36 98.68 57.28 40.14 51.71 98.89 67.91 51.41
SalsaNext[32] 73.12 80.48 75.23 77.76 63.21 93.79 84.21 88.74 79.76 82.74 90.89 86.62 76.40
Cylinder3D[33] 95.41 98.85 95.56 97.18 94.51 98.94 96.37 97.64 95.38 99.24 97.06 98.14 96.34
WeatherNet[20] 71.73 72.96 85.40 78.69 64.87 84.97 80.54 82.69 70.49 80.44 99.06 88.78 79.83
4DenoiseNet[21] 84.38 78.97 96.86 87.01 76.99 92.34 97.53 94.86 90.23 87.56 97.88 92.43 85.93
3D-OutDet[22] 90.87 95.17 90.43 92.74 86.46 97.53 90.16 93.70 88.15 99.44 98.53 98.98 97.99
TripleMixer (Ours) 97.21 97.32 98.37 97.84 95.78 99.37 98.02 98.69 97.42 99.72 98.71 99.21 98.43
Figure 7: Visualization of Denoising Performance Results on Weather-KITTI and Weather-NuScenes Datasets. Our TripleMixer demonstrates consistently stable performance across different scenarios.
Figure 8: Qualitative results among the proposed TripleMixer and state-of-the-art denoising models[21, 22, 33] on the WADS[17], the proposed Weather-KITTI and Weather-NuScenes datasets. To highlight the differences, the correct / incorrect predictions are painted in gray / red, respectively. Best viewed in colors and zoomed-in for details.
Results On Weather-KITTI: In the previous section, we demonstrated the superior performance of our denoising model on the real-world WADS[17] dataset. To further validate the good generalizability of our model, we conducted additional denoising comparison experiments on our proposed Weather-KITTI and Weather-NuScenes datasets. For a thorough evaluation, we selected the most recent and best-performing models from Table LABEL:wads for further comparative analysis.
As shown in Table LABEL:Weather-KITTI, we compared our model with six state-of-the-art methods across three scenarios in the Weather-KITTI dataset: Snow Scenarios, Fog Scenarios and Rain Scenarios. Our model achieved the highest average mIoU of 96.31 across these three scenarios, surpassing popular general semantic segmentation methods such as Cylinder3D[33], as well as the latest denoising models like 4DenoiseNet[21] and 3D-OutDet[22]. Specifically, our model’s average mIoU on the Weather-KITTI dataset is 2.2% higher than Cylinder3D[33], 16.2% higher than 4DenoiseNet[21], and 5.9% higher than 3D-OutDet[22]. It is noteworthy that our model consistently achieved the highest mIoUs of 96.46, 95.61, and 96.87 across Snow, Fog, and Rain scenarios, respectively, while DROR[30], Cylinder3D[33], and SalsaNext[32] achieved the highest Recall or Precision in the Snow, Fog, and Rain scenarios, respectively. This demonstrates the robustness of our model in point cloud denoising on the Weather-KITTI dataset.
Results On Weather-NuScenes: We also compared the denoising performance of our model with six state-of-the-art methods across three scenarios in the Weather-NuScenes dataset. As shown in Table LABEL:Weather-NuScenes, our TripleMixer model achieved the highest average mIoU of 97.21 across these three scenarios. Notably, the average mIoU of our model is 6.9% higher than 3D-OutDet[22] and 15.20% higher than 4DenoiseNet[21]. Additionally, our model achieved the best mIoU accuracy in each scenario (Snow, Fog, and Rain) compared to other denoising models, with mIoUs of 95.78, 97.42, and 98.48, respectively. These results demonstrate that our model exhibits superior performance in the Weather-NuScenes dataset across three different adverse weather scenarios compared to other denoising models.
We further present the mIoU performance of our model and three other denoising models across six scenarios in the Weather-KITTI and Weather-NuScenes datasets, as shown in Figure 7. Our model not only achieved the highest mIoU but also exhibited the smallest variance in mIoU accuracy across the six scenarios, with a variance of only 0.96. This indicates that our denoising model has strong generalization capability and is the most robust compared to the other denoising models, effectively denoising across diverse weather conditions.
Visualization: We visualized our model’s prediction results for point cloud denoising performance on WADS[17], Weather-KITTI, and Weather-NuScenes, compared with those of the state-of-the-art models, including: 4DenoiseNet[21], 3D-OutDet[22], and Cylinder3D[33]. As illustrated in Figure 8, it is evident that our TripleMixer model delivers more accurate predictions, achieving optimal denoising results across different weather scenarios compared to the other three denoising models.
VI Discussions
VI-A Ablation Study:
In this section, we perform thorough ablation studies on the point cloud denoising task to investigate the effect of each components, projection plane resolution, wavelet decomposition levels and loss functions of our TripleMixer model. All ablation experiments are conducted on the real-world WADS[17] dataset, with same experimental settings as in Section V.
Table VII: Ablation Results of Each Component. GMX: Geometry Mixer Layer, FMX: Frequency Mixer Layer, and CMX: Channel Mixer Layer.
Experiment GMX FMX CMX Metric(%)
Precision Recall F1 mIOU
I 95.44 93.86 94.64 89.83
II 95.11 93.65 94.37 89.35
III 96.34 93.71 95.01 90.48
IV 96.38 93.93 95.13 90.73
Table VIII: Ablation Results of Different XY Directions Projection Resolution When Z Direction Projection Resolution is set to 32
XY-Resolution Precision↑ Recall↑ F1↑ mIoU↑
[64,64] 96.24 93.63 94.91 90.32
[128,128] 96.55 93.57 95.03 90.54
[256,256] 96.38 93.93 95.13 90.73
[384,384] 96.33 98.83 95.06 90.59
[512,512] 96.25 98.85 95.03 90.53
Table IX: Ablation Results of Different Z Direction Projection Resolution When XY Directions Projection Resolution is set to [256,256]
Z-Resolution Precision↑ Recall↑ F1↑ mIoU↑
[16] 96.43 93.73 95.06 90.57
[24] 96.55 93.70 95.10 90.66
[32] 96.38 93.93 95.13 90.73
[40] 96.43 93.69 95.04 90.54
[48] 96.46 93.59 95.00 90.48
Table X: Ablation Results of Different Wavelet Decomposition Levels in the Frequency Mixer Layer
Methods Precision↑ Recall↑ F1↑ mIoU↑
w/o Wavelet 95.11 93.65 94.37 89.35
w/ Wavelet-1 96.33 93.54 94.91 90.32
w/ Wavelet-2 96.38 93.93 95.13 90.73
w/ Wavelet-3 96.16 93.80 94.96 90.41
[…]
F1 mIOU
I 94.84 90.19
II 95.05 90.56
III 94.98 90.45
IV 95.13 90.73
Table XII: Semantic Segmentation Performance Comparison on WADS Datasets. The Row Marked with $\Delta$ Represents the Performance Difference of Our Denoising Model TripleMixer for Downstream Semantic Segmentation Tasks.
Methods mIoU
car
truck
other-vehicle
person
road
parking
sidewalk
other-ground
building
fence
other-structure
vegetation
trunk
terrain
pole
traffic-sign
other-objec
acc-snow
SalsaNext[32] 14.32 45.98 4.84 0.00 0.00 50.56 4.21 6.53 2.47 46.85 2.40 1.05 47.35 0.34 0.31 6.12 7.46 0.29 29.91
SalsaNext[32] + TripleMixer 15.61 50.22 1.20 0.45 0.93 53.83 14.67 8.79 7.92 54.97 3.86 0.50 43.25 0.29 0.35 7.08 11.00 1.41 20.22
$\Delta$ +1.29 +4.24 -3.64 +0.45 +0.93 +3.27 +10.46 +2.26 +5.45 +8.12 +1.46 -0.55 -4.10 -0.05 +0.04 +0.96 +3.54 +1.12 -9.69
Cylinder3D[33] 19.40 61.88 0.00 0.00 0.00 47.45 0.15 4.46 0.17 72.01 16.03 1.01 51.76 0.00 0.00 24.40 14.42 0.36 54.16
Cylinder3D[33] + TripleMixer 21.03 66.03 0.00 0.26 0.48 46.90 2.20 5.99 0.57 79.13 14.47 1.37 57.80 0.18 0.24 31.19 17.92 1.62 52.28
$\Delta$ +1.63 +4.15 0.00 +0.26 +0.48 -0.55 +1.05 +1.53 +0.40 +7.12 -1.56 +0.36 +6.04 +0.18 +0.24 +6.79 +3.50 +1.26 -1.88
PolarNet[50] 19.28 59.10 0.03 0.00 0.00 48.25 0.01 0.14 0.20 72.30 3.84 0.01 61.94 0.00 0.00 19.94 21.77 0.39 59.27
PolarNet[50] + TripleMixer 20.36 60.21 0.08 0.12 0.06 59.38 0.01 0.01 0.04 68.40 13.29 0.42 59.08 0.22 0.26 27.45 21.78 2.32 53.23
$\Delta$ +1.08 +1.11 +0.05 +0.12 +0.06 +11.13 0.00 -0.13 -0.16 -3.90 +9.45 +0.41 -2.86 +0.22 +0.26 +7.51 +0.01 +1.93 -6.04
Effects of Each Component. To testify the effectiveness of each component of TripleMixer, we conduct an extensive ablation study and list the result in Table VII. As shown in Table VII, Experiments I-III represent the accuracy of our TripleMixer model after sequentially removing the GMX Layer, the FMX Layer, and the CMX Layer, respectively. From Experiments I and IV, we observe that the GMX Layer contributes significantly, resulting in F1 and mIoU scores of 0.49 and 0.90, respectively. Similarly, Experiments II and IV indicate that the FMX Layer enhances performance, yielding F1 and mIoU improvements of 0.65 and 1.19, respectively. Finally, the impact of the CMX Layer is evident from Experiments III and IV, showing F1 and mIoU increases of 0.12 and 0.25, respectively. These results demonstrate that the GMX Layer, the FMX Layer, and the CMX Layer are all crucial for improving model performance, each contributing significantly to the overall effectiveness of our TripleMixer model.
Effects of Projection Plane Resolution. In the Frequency Mixer Layer, the 3D geometric features are projected onto three 2D planes along the X, Y, and Z axes. An adaptive wavelet decomposition is then performed on these projected 2D planes to extract multi-scale features. The initial projection resolution of each 2D plane influences the effectiveness of the multi-scale feature extraction. In this part, we conduct ablation experiments to assess the impact of different projection plane resolutions on the performance of our TripleMixer model. Due to the different horizontal and vertical emission ranges of LiDAR, we set different projection resolutions for the XY directions and the Z direction. We first set the projection resolution for the Z direction to 32 in order to select the optimal projection resolution for the XY directions. Table LABEL:plane_reso_xy shows the impact of different projection resolutions in the XY directions on the model’s denoising performance, with the Z direction projection resolution set to 32. As shown in Table LABEL:plane_reso_xy, the optimal mIoU accuracy is achieved when the projection resolution in the XY directions is set to [256,256]. Both excessively high resolutions such as [512,512] and excessively low resolutions such as [64,64] result in suboptimal performance. Therefore, we selected the resolution of [256,256] for the XY directions. We further verified the impact of different resolutions in the Z direction on model accuracy. The results are shown in Table LABEL:plane_reso_z, where we experimented with five different resolutions for the Z direction. As indicated in Table LABEL:plane_reso_z, with the XY directions’ resolution fixed, the optimal mIoU accuracy is achieved when the Z direction resolution is set to 32. Similar to the results for the XY directions, excessively high or low resolutions in the Z direction also resulted in poorer accuracy. Therefore, in our practical experiments, we chose the projection resolution of [256,256,32] for the XYZ directions.
Effects of Wavelet Decomposition Levels. In this part, we investigate the impact of varying levels of wavelet decomposition in the Frequency Mixer Layer on our model’s performance. With the projection resolution set to [256,256,32] for the XYZ directions, we select a maximum wavelet decomposition level of 3. This choice is due to the fact that if the wavelet decomposition level exceeds 3, the resolution in the Z direction becomes $\leq 2$, which hinders the extraction of effective information. Consequently, we analyze the model’s performance across a range of wavelet decomposition levels, from none to the maximum level of 3. As shown in Table LABEL:wavelet, not employing any wavelet decomposition levels results in the poorest model accuracy, highlighting the critical role that wavelet transformations play in optimizing model performance. Additionally, utilizing the maximum number of decomposition levels also leads to suboptimal performance, as excessive wavelet decomposition fails to extract more effective information. Our TripleMixer achieves the best overall accuracy when the decomposition level is set to 2.
Effects of Loss Functions. Our model’s loss function primarily consists of three components: the cross-entropy loss, the Lovász loss, and the wavelet regularization terms. We use the cross-entropy loss as the baseline to test the impact of different loss functions on model accuracy. As shown in Table XI, after adding only the Lovász loss, the mIoU and F1 scores of our TripleMixer model improved by 0.37 and 0.21, respectively. Additionally, after incorporating only the wavelet regularization terms, the model’s mIoU and F1 scores increased by 0.26 and 0.14, respectively. When using the combination of all three loss functions, our denoising model achieved the highest mIoU and F1 scores.
VI-B Experiments on Segmentation Tasks
In this section, we apply the proposed denoising model to LiDAR-based downstream perception tasks. We conducted point cloud semantic segmentation experiments on the WADS dataset, both with and without our denoising model. These experimental results demonstrate the generalization ability of our proposed method and its effectiveness on LiDAR point cloud processing.
The performance of the point cloud semantic segmentation task for SalsaNext[32], Cylinder3D[33] and PolarNet[50] models before and after removing snow with our TripleMixer model is shown in Table LABEL:wads_denoise, where $\Delta$ indicates the performance differences resulting from the snow removal. As illustrated in Table LABEL:wads_denoise, with the assistance of our model, the semantic segmentation performance of the SalsaNext[32], Cylinder3D[33], and PolarNet[50] models has significantly improved. The performance enhancements are quantified as increases of 1.29%, 1.63%, and 1.08%, respectively. Notably, before the snow removal by TripleMixer, none of the three models were capable of detecting the ’other-vehicle’ and ’person’ categories. However, with the integration of our model, there is a marked improvement in the detection accuracy for these categories. This demonstrates that our model is beneficial and versatile for LiDAR-based downstream perception tasks.
VII Conclusion and Future Work
In this work, we first introduce two large-scale adverse weather point cloud datasets, Weather-KITTI and Weather-NuScenes, featuring a wide range of weather scenarios, complete original LiDAR information, and point-wise weather and semantic labels. These datasets provide robust data support for further research into adverse weather conditions. Subsequently, we propose a novel point cloud denoising model, the TripleMixer network, which integrates spatial geometric features, frequency domain features, and multi-channel contextual information to achieve high denoising accuracy. This model demonstrates optimal performance on both real-world datasets and our newly proposed datasets. Finally, we validate the significant performance enhancement of our denoising model for downstream point cloud tasks, specifically semantic segmentation, under adverse weather conditions. In future work, we aim to incorporate additional sensors, such as cameras and IMUs, to further explore the perception capabilities of multi-sensor fusion in extreme weather conditions.

### v2

@@
 Abstract
-LiDAR sensors are crucial for providing high-resolution 3D point cloud data in autonomous driving systems, enabling precise environmental perception.
-However, real-world adverse weather conditions, such as rain, fog, and snow, introduce significant noise and interference, degrading the reliability of LiDAR data and the performance of downstream tasks like semantic segmentation.
-Existing datasets often suffer from limited weather diversity and small dataset sizes, which restrict their effectiveness in training robust models.
-Additionally, current deep learning denoising methods, while effective in certain scenarios, often lack interpretability, complicating the ability to understand and validate their decision-making processes.
-To overcome these limitations, we introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions: rain, fog, and snow.
-These datasets retain the original LiDAR acquisition information and provide point-level semantic labels for rain, fog, and snow.
-Furthermore, we propose a novel point cloud denoising model, TripleMixer, comprising three mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer.
-These layers are designed to capture geometric spatial information, extract multi-scale frequency information, and enhance the multi-channel feature information of point clouds, respectively.
-By leveraging the combined capabilities of these three core mixer layers, the TripleMixer network not only effectively preserves and enhances critical features of point cloud data during the denoising process but also offers strong interpretability.
-Experiments conducted on the WADS dataset in real-world scenarios, as well as on our proposed Weather-KITTI and Weather-NuScenes datasets, demonstrate that our model achieves state-of-the-art denoising performance.
-Additionally, our experiments show that integrating the denoising model into existing segmentation frameworks enhances the performance of downstream tasks.
-These results highlight our model’s potential to significantly improve the accuracy and reliability of autonomous driving systems in complex environments.
-The datasets and code will be made publicly available at https://github.com/Grandzxw/TripleMixer.
+Adverse weather conditions such as snow, fog, and rain pose significant challenges to LiDAR-based perception models by introducing noise and corrupting point cloud measurements.
+To address this issue, we propose TripleMixer, a robust and efficient point cloud denoising network that integrates spatial, frequency, and channel-wise processing through three specialized mixer modules.
+TripleMixer effectively suppresses high-frequency noise while preserving essential geometric structures and can be seamlessly deployed as a plug-and-play module within existing LiDAR perception pipelines.
+To support the development and evaluation of denoising methods, we construct two large-scale simulated datasets, Weather-KITTI and Weather-NuScenes, covering diverse weather scenarios with dense point-wise semantic and noise annotations.
+Based on these datasets, we establish four benchmarks: Denoising, Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD).
+These benchmarks enable systematic evaluation of denoising generalization, transferability, and downstream impact under both simulated and real-world adverse weather conditions.
+Extensive experiments demonstrate that TripleMixer achieves state-of-the-art denoising performance and yields substantial improvements across all downstream tasks without requiring retraining.
+Our results highlight the potential of denoising as a task-agnostic preprocessing strategy to enhance LiDAR robustness in real-world autonomous driving applications.
 Index Terms:
-LiDAR Dataset, Weather-KITTI, Weather-Nuscenes, Point Cloud Denosing, Adverse Weather
+Adverse Weather Dataset, Point Cloud Denosing, Robust LiDAR perception, Plug-and-Play Network
 I Introduction
-Figure 1: Visualization of semantic 3D points from the same LiDAR frame across four weather conditions in our proposed Weather-KITTI dataset. Weather noises are marked in red.
-LiDAR sensors play a crucial role in autonomous driving systems by emitting laser beams and capturing the reflected signals to generate high-resolution 3D point cloud data, thus providing precise environmental perception[1, 2, 3, 4].
-However, the LiDAR data included in current large-scale autonomous driving datasets are typically ”clean”.
-In real-world scenarios, varying weather conditions, such as rain, fog, and snow, can cause laser beams to scatter, be absorbed, or reflect abnormally, leading to significant noise and interference[5, 6, 7].
-This increase in noise not only reduces the reliability of the LiDAR data but also potentially degrades the accuracy of downstream tasks, such as semantic segmentation and object detection, which in turn can affect the overall decision-making capability of autonomous driving systems[8, 9, 10, 11].
-Therefore, effectively denoising LiDAR data while preserving its accuracy under diverse weather conditions has become a critical challenge in autonomous driving systems.
-To address this challenge, some researchers have specifically developed LIDAR datasets under adverse weather conditions.
-These datasets can be broadly categorized into two types: real-world datasets, collected under adverse weather conditions, and synthetic datasets, generated through computer simulations[12, 13, 14].
-Real-world datasets often face challenges related to the complexity of weather annotation, high data collection costs, and small dataset sizes, which restrict their applicability across a wide range of weather conditions.
-On the other hand, synthetic datasets allow researchers to create extreme weather scenarios under controlled conditions, enabling more comprehensive testing of system reliability.
-However, existing synthetic datasets suffer from limitations such as limited weather types, the loss of original LiDAR acquisition information and a lack of precise point-level weather labels.
-In addition to these datasets, several denoising methods for point clouds have been proposed, which can be broadly categorized into statistical approaches and deep learning methods[12, 13].
-Statistical approaches filter noise by analyzing the distribution patterns of point clouds.
-However, despite their effectiveness on smaller datasets, they often require substantial computational resources, especially when dealing with large-scale point clouds[15, 16, 17, 18, 19]; In contrast, deep learning methods leverage the powerful feature-learning capabilities of neural networks to directly learn noise patterns from input point clouds.
-Existing deep learning denoising networks are typically further classified into two categories based on the input format of the point clouds: range image-based methods and 3D point cloud-based methods[20, 21, 22, 23].
-The former involves projecting 3D point cloud data onto a 2D range images, but this projection process may lead to a loss of spatial information, potentially affecting the final denoising performance, particularly when dealing with complex point cloud data.
-The latter directly processes raw 3D point cloud data, better preserving the spatial structure and geometric information, which results in more effective denoising.
-However, while these deep learning methods have achieved promising results in denoising tasks, they often lack interpretability.
-To address the challenges posed by existing datasets and denoising methods, this paper presents two key solutions.
-First, we introduce two large-scale adverse weather datasets, Weather-KITTI and Weather-NuScenes, which cover three common adverse weather conditions, including rain, fog, and snow.
-These datasets retain the original LiDAR acquisition information, such as intensity, and provide point-level labels for rain, fog, and snow.
-Our datasets not only better reflect the performance of LiDAR sensors under different weather conditions in the real world but also provide researchers with more precise and detailed annotations, supporting deeper algorithm research and model optimization.
-This design aims to provide robust and reliable data support for improving the robustness and reliability of autonomous driving systems under complex weather conditions.
-Second, we propose a novel point cloud denoising model, the TripleMixer network, which consists of three mixer layers: the Geometry Mixer Layer, Frequency Mixer Layer, and Channel Mixer Layer.
-The Geometry Mixer Layer focuses on capturing the geometric spatial information of point clouds by analyzing and mixing the geometric relationships of neighboring points, thereby preserving local structural features and important geometric details during the denoising process.
-The Frequency Mixer Layer is dedicated to extracting frequency information from point clouds by employing multi-scale analysis.
-This process effectively learns the distinct spectral features from low-frequency components, which represent the global structure, and from high-frequency components, which contain both details and noise, improving the accuracy of point cloud interpretation in complex environments.
-The Channel Mixer Layer enhances the multi-channel feature information of point clouds by facilitating information exchange and integration between different feature channels, capturing contextual relationships between features, and further improving denoising accuracy.
-The collaborative effect of these three core mixer layers enables the TripleMixer network to not only preserve and enhance critical features of point cloud data during the denoising process but also provides strong interpretability.
-Experiments conducted on the WADS dataset[17] in real-world scenarios, as well as on our proposed weather-KITTI and Weather-Nuscene datasets, demonstrate that our model achieves state-of-the-art denoising performance.
-Additionally, when applying our denoising model to downstream semantic segmentation tasks, experimental results indicate that integrating our denoising model into existing segmentation models can notably improve their segmentation performance.
+Figure 1: Visualization of semantic 3D points from the same LiDAR frame across four weather conditions in our proposed Weather-KITTI dataset.
+Weather noises are marked in red and loss points are marked in pink.
+Robust LiDAR perception [1, 2] in adverse weather remains a critical challenge for autonomous driving systems.
+Weather-induced degradations such as snow, fog, and rain can cause LiDAR beams to scatter, reflect, or be absorbed abnormally, leading to corrupted point cloud measurements with substantial noise and missing points [3].
+These degradations compromise the reliability of 3D sensing and degrade the performance of downstream perception tasks [4, 5, 6, 7, 8], including semantic segmentation, place recognition, and object detection.
+To address this issue, recent research has explored two broad categories of solutions: domain transfer methods [9, 10, 11] and data preprocessing via denoising [12, 13, 14].
+Domain transfer methods, including domain adaptation and domain generalization techniques, aim to improve model robustness by adapting features across clean and adverse-weather domains.
+However, they typically require access to target domain data, involve task-specific retraining, and often lack generalizability across different perception tasks.
+In contrast, denoising serves as a task-agnostic preprocessing step that can be seamlessly integrated into existing LiDAR perception pipelines without the need for retraining.
+Existing denoising approaches can be broadly categorized into statistical and learning-based methods [15, 16].
+Statistical methods [17, 18, 19] rely on local geometric heuristics, but their iterative filtering procedures make them computationally inefficient and poorly scalable to large-scale point clouds.
+Learning-based methods typically follow one of two paradigms: projecting 3D point clouds onto 2D range images [20, 21] or directly processing raw point clouds [14].
+Most existing models exhibit limited ability to capture multi‑scale spatial structural information and lack frequency‑aware modeling to distinguish meaningful structures from high‑frequency noise, resulting in suboptimal denoising performance under adverse weather conditions.
+To overcome these challenges, we propose TripleMixer, a robust and efficient point cloud denoising architecture tailored for adverse weather conditions.
+TripleMixer consists of three complementary components: (1) a Geometry Mixer (GMX) that captures local geometric structures from raw 3D points while preserving fine-grained spatial details; (2) a Frequency Mixer (FMX) that performs multi-scale frequency decomposition across three orthogonal 2D projections to effectively separate structural information from high-frequency noise; and (3) a Channel Mixer (CMX) that facilitates inter-channel feature refinement to enhance contextual representation.
+This design enables TripleMixer to suppress noise while preserving critical geometric features, striking an effective balance between performance, efficiency, and interpretability.
+To advance the research community’s capabilities in robust LiDAR perception under adverse weather, and to support the development and rigorous evaluation of TripleMixer, we construct two large-scale simulated datasets: Weather-KITTI and Weather-NuScenes.
+As shown in Table I, our datasets contain over 200K LiDAR frames, substantially exceeding the scale of existing benchmarks, which typically include fewer than 50K frames.
+Moreover, our datasets preserve the characteristics of real LiDAR acquisitions while covering multiple adverse weather types (snow, fog, and rain).
+They also provide both point-wise semantic labels and weather annotations, which are lacking in most existing datasets.
+Built upon these datasets, we establish four benchmarks: (1) Denoising benchmark for evaluating denoising performance in isolation, and three downstream perception benchmarks under real-world adverse weather conditions: (2) Semantic Segmentation (SS), (3) Place Recognition (PR), and (4) Object Detection (OD).
 Our main contributions are summarized as follows:
-  • We introduce two large-scale datasets, Weather-KITTI and Weather-NuScenes, specifically designed for autonomous driving in three adverse weather conditions: rain, fog, and snow.
-These datasets retain the original LiDAR acquisition information, including intensity, and provide precise point-level weather labels.
-We believe that these datasets will support efforts to improve the robustness and reliability of autonomous driving systems under complex weather conditions.
-  • We propose a novel point cloud denoising model, the TripleMixer network, which consists of three distinct mixer layers: the Geometry Mixer Layer, the Frequency Mixer Layer, and the Channel Mixer Layer.
-These layers effectively preserve geometric spatial information, separate frequency components, and enhance multi-channel feature integration.
-Our TripleMixer network not only preserves and enhances critical features of point cloud data during the denoising process but also provides strong interpretability.
-  • Extensive experiments conducted on both existing and newly proposed datasets demonstrate the state-of-the-art denoising performance of the proposed TripleMixer model, and show that integrating our denoising model into existing semantic segmentation frameworks leads to notable improvements in downstream tasks.
-These findings highlight the practical applicability and effectiveness of the proposed model in enhancing the robustness and reliability of autonomous driving systems in real-world scenarios.
+  • We propose TripleMixer, a plug-and-play point cloud denoising network that integrates spatial, frequency, and channel-wise processing through three specialized mixer layers.
+TripleMixer enables interpretable and robust denoising under adverse weather conditions, and can be seamlessly integrated into existing LiDAR perception pipelines to enhance their robustness.
+  • We introduce Weather KITTI and Weather NuScenes, two large-scale simulated LiDAR datasets encompassing diverse adverse weather conditions.
+Based on these datasets, we establish four benchmarks: Denoising, SS, PR, and OD.
+These benchmarks support comprehensive evaluation and provide a solid foundation for developing robust denoising and 3D perception models under both simulated and real-world adverse weather conditions.
+  • Extensive experiments on both all benchmarks demonstrate that TripleMixer achieves state-of-the-art denoising performance.
+Furthermore, when used as a preprocessing module without retraining downstream models, TripleMixer significantly improves performance on SS, PR, and OD tasks under real-world adverse weather, achieving relative accuracy gains of up to 20%, 56%, and 16%, respectively.
 II Related Work
-II-A Adverse Weather Datasets
-Weather changes are common in the physical world of autonomous driving.
-Recent datasets investigating LiDAR performance in adverse weather conditions are generally divided into two categories: real-world recordings and synthetic point clouds.
-Existing real-world datasets face challenges with complex manual weather labeling and limited scale.
-Meanwhile, current synthetic datasets often fail to retain the original characteristics of LiDAR.
-Real-world datasets are typically collected either outdoors or in weather chambers.
-The DENSE dataset[24] was recorded within fog chambers, which allow for full control over weather conditions and surroundings.
-However, despite this level of control, these chambers do not adequately reflect real-world conditions.
-The CADC dataset[25] was recorded during the wintertime in adverse Canadian weather with the VLP32 LiDAR sensors.
-The dataset has 3D bounding box annotations available, but does not come with weather-related labels and it only covers snowfall weather conditions.
-The Boreas dataset[26] was collected by driving a repeated route under adverse weather conditions such as rain and falling snow, using a 128-channel LiDAR.
-It provides 3D bounding box annotations for 3D object detection, but also does not offer weather-related labels.
-The WADS dataset[17] is an extreme weather dataset from open-world scenarios and includes point-wise weather labels.
-However, due to the complexity of manually labeling snow noise points individually, it features a limited scale and few scenarios.
-Synthetic datasets are collected by augmenting existing datasets.
-The SnowKITTI dataset[21] is a synthetic dataset based on the KITTI odometry benchmark dataset[27].
-However, it only contains data from snowfall weather conditions and has altered the original KITTI point cloud format, with each point cloud frame formatted as 1048×64.
-Additionally, it has lost the intensity information of each point cloud.
-The SMART-Degradation dataset[28] provides additional raw data collected by a 32-channel 3D LiDAR and synthesized rainy LiDAR scans.
-However, it does not provide point-wise weather labels for the rainy LiDAR scans, making it challenging to apply to point cloud perception tasks.
-II-B Point Cloud Denoising Method
-Adverse weather conditions can impair the effectiveness of LiDAR sensors by introducing noises into their measurements.
-These noises can significantly degrade the performance of tasks such as perception and mapping[29].
-Consequently, denoising point clouds under such conditions is essential to restore measurement clarity.
-The emerging field of denoising LiDAR point clouds in adverse weather conditions has been explored in only a few studies so far.
-The primary methodologies employed include statistical approaches and deep learning approaches.
-Statistical approaches filter noise by analyzing the distribution patterns of point clouds.
-Statistical Outlier Removal (SOR) and Radius outlier removal (ROR) [16] filter out noise by analyzing the density of the nearest neighbor points.
-SOR removes a noisy point by comparing the mean and standard deviation of the distance to its k-nearest neighbors.
-ROR processes each point in the point cloud, counting the number of neighbors within a specified search radius to remove noise points.
-Dynamic ROR (DROR) [30] enhances the ROR filter by dynamically adjusting the search radius.
-Dynamic SOR (DSOR) [17] is an extension of SOR filter designed to address the non-uniform density within point clouds.
-It dynamically calculates a local distance threshold to remove noise points.
-Low-Intensity Outlier Removal (LIOR)[19] and Dynamic Distance–Intensity Outlier Removal (DDIOR)[18] integrates the distance and intensity of points based on the systematic and accurate analysis of data characteristics in snowy weather.
-Conditional Random Fields Outlier Removal (CRFOR)[31] proposes a de-snowing formulation with Conditional Random Fields and Spatio-Temporal information.
-Since statistical approaches typically remove noise through iterative algorithms, they are computationally expensive.
-Contrary to statistical approaches, deep learning approaches became popular due to the model’s abilities to directly understand the characteristics of weather-induced noise in point cloud.
-Some general-purpose point cloud semantic segmentation approaches [32, 33, 34] have been adapted to denoise point clouds affected by adverse weather.
-Although they have achieved good performance, they are not specifically designed for denoising.
-Existing denoising networks are typically categorized based on the input format of the point cloud: range images and raw 3D point clouds.
-WeatherNet[20] uses modified LiLaBlock[35] to filter out adverse weather effects in point cloud data.
-4DenoiseNet[21] proposes a k-nearest neighbor convolution kernel that captures spatial-temporal information, and with a attention mechanism.
-However, its limitation lies in the requirement for LiDAR scans with temporal continuity.
-LiSnowNet[36] combines the Fast Fourier Transform (FFT) with Convolutional Neural Networks (CNNs) for unsupervised de-noising to remove snow from LiDAR point clouds.
-The denoising networks mentioned above all use range projection images as inputs.
-While range projection methods usually benefit from fast computations, they face limitations such as projection errors and loss of spatial information when converting 3D point clouds into 2D range images.
-3D-OutDet[22] takes raw 3D point clouds as inputs and develops a NeighborHood (NH) Convolution for unorganized scattered 3D point clouds, specifically targeting the removal of snowflakes.
-Table I: Comparison of the Proposed Adverse Weather Dataset with Existing Public Datasets.
-Including: Year, LiDAR Parameters, Number of Dataset Frames, Presence of Semantic Labels, Types of adverse Weather Covered,
-Inclusion of LiDAR Intensity Features, and Presence of Point-Wise Weather Labels.
+II-A Robust LiDAR Perception
+Adverse weather conditions pose significant challenges to LiDAR-based perception by introducing severe noise and degradation into point cloud measurements [22].
+Recent efforts to address this issue primarily fall into two categories: domain transfer and data preprocessing through denoising.
+Domain transfer methods include domain adaptation [9, 23, 11], which leverages labeled or unlabeled target-domain (i.e., adverse weather) data to align feature distributions via retraining or fine-tuning, and domain generalization [24, 10], which enhances model robustness through data augmentation or domain-invariant representation learning.
+However, these approaches typically require access to target-domain data, involve task-specific retraining, and often struggle to generalize across diverse perception tasks.
+In contrast, denoising offers a task-agnostic and modular preprocessing solution that can be seamlessly integrated into existing LiDAR perception pipelines.
+Existing denoising approaches can be broadly categorized into statistical and deep learning-based methods.
+Statistical methods, such as SOR and ROR [17], rely on local geometric heuristics, while dynamic variants like DROR [25] and DSOR [18], as well as intensity-aware filters such as LIOR [19] and DDIOR [26], offer improved resilience to weather-induced noise.
+However, their iterative nature makes them computationally expensive and limits scalability.
+Deep learning-based approaches learn complex noise patterns from data and are typically grouped by input modality: 2D range images or raw 3D point clouds.
+Range image-based methods, such as WeatherNet [20], LiSnowNet [27], and 4DenoiseNet [21], leverage CNNs or FFTs to process projected range representations, but inherently compromise the fidelity of 3D spatial structure.
+In contrast, point cloud-based models like 3D-OutDet [14] directly operate on raw point clouds using neighborhood-based convolutions, preserving geometric detail but introducing significant runtime overhead due to $k$-NN preprocessing.
+Despite recent progress, existing methods remain limited in their capacity to capture multi-scale spatial structure and generally lack frequency-aware modeling, which is critical for distinguishing meaningful features from high-frequency noise.
+II-B Adverse Weather Datasets and Benchmarks
+Existing adverse-weather datasets for LiDAR perception can be broadly categorized into two types: real-world datasets and simulated datasets.
+Real-world datasets are collected using actual sensors under varying environmental conditions, providing realistic examples of sensor behavior in adverse weather.
+For instance, the STF dataset [28] offers multimodal recordings with different levels of visibility, while the CADC dataset [29] captures winter driving scenes using a VLP-32 LiDAR and provides 3D bounding box annotations.
+The WADS dataset [18] focuses on extreme outdoor environments and includes point-wise annotations of weather-induced noise, making it suitable for fine-grained analysis.
+Simulated datasets are typically created by applying weather simulators to existing clean-weather LiDAR datasets.
+These models simulate the physical interactions between LiDAR beams and environmental factors such as rain, fog, and snow.
+For example, a hybrid Monte Carlo model is employed in [30] to simulate raindrop interference, while fog effects are modeled using data-driven and physically based methods in [31, 32].
+Snow simulation is handled in [33] by superimposing reflections from virtual particles.
+Building on these techniques, datasets such as SnowKITTI [21] and Robo3D [34] augment existing datasets like KITTI [35], SemanticKITTI [36], and nuScenes [37] with weather effects.
+Some recent benchmarks [38, 39] also introduce synthetic perturbations across multiple sensor modalities to evaluate model robustness.
+Overall, real-world datasets offer valuable realism by capturing authentic environmental variations and sensor responses.
+However, their scale is often constrained by the high cost of manual annotation, and they typically provide limited coverage across diverse weather conditions.
+In contrast, simulated datasets enable controllable generation of adverse weather effects and can effectively augment existing clean-weather datasets.
+Nonetheless, most existing simulated datasets are restricted in scale, lack fine-grained weather annotations, and have yet to be systematically validated for their effectiveness in enhancing downstream perception tasks under real‑world adverse weather conditions.
+To address these limitations, we introduce two large-scale simulated datasets and establish four benchmarks spanning both denoising and downstream LiDAR perception tasks under real-world adverse weather scenarios.
+Table I: Comparison of the Proposed Adverse Weather Dataset with Existing Public Datasets, including Year, LiDAR Parameters, Frame Count, Semantic Labels, Weather Types, Intensity Features, and Point-Wise Weather Noise Labels.
 Dataset Year LiDAR Size Semantic Labels Weather Types Intensity Weather Labels
@@
 WADS[18] 2021 64 channel 1K frames point-wise
-CADC[25] 2021 32 channel 7K frames bounding boxes
-DENSE[24] 2020 32 & 64 channel 13K frames bounding boxes
-Boreas[26] 2023 128 channel 7K frames bounding boxes
-SMART-Degradation[28] 2021 32 channel 1K frames -
-Oxford RobotCar[37] 2020 32 channel 11K frames -
-SemanticKITTI[38] 2019 64 channel 2K frames point-wise
-nuScenes-lidarseg[39] 2020 32 channel 28K frames point-wise
-SnowKITTI[21] 2023 64 channel 43K frames -
+CADC[29] 2021 32 channel 7K frames object-wise
+STF[28] 2020 32 & 64 channel 13K frames object-wise
+LIBRE[40] 2020 32 & 64 & 128 channel 6K frames object-wise
+Boreas[41] 2023 128 channel 7K frames object-wise
+SemanticSTF[23] 2023 64 channel 2K frames point-wise
+Robo3d-KITTI[34] 2023 64 channel 8K frames object-wise
+Robo3d-SemanticKITTI[34] 2023 64 channel 8K frames point-wise
+Robo3d-nuScenes[34] 2023 32 channel 12K frames point-wise
+Robo3d-WOD[34] 2023 64 channel - - -
+SnowKITTI[21] 2023 64 channel 43K frames
 Weather-KITTI (Ours) 2024 64 channel 130K frames point-wise
 Weather-NuScenes (Ours) 2024 32 channel 84K frames point-wise
-III Datasets Generation
-As introduced in the previous section, current adverse weather datasets face several challenges: Firstly, there is an absence of point-wise semantic labels and weather labels; Secondly, manually annotating weather conditions in large-scale, real-world datasets is highly complex; Thirdly, existing synthetic datasets lack the original LiDAR acquisition information such as intensity, which impacts the performance of LiDAR downstream perception tasks.
-Building on this, we propose our synthetic adverse weather datasets, named Weather-KITTI and Weather-NuScenes, which are based on the SemanticKITTI[38] and nuScenes-lidarseg[39] datasets, respectively.
+  • : SemanticSTF provides weather‑related labels for indiscernible points under adverse weather but lacks separate noise annotations,
+    ✓: included, ✗: not included, –: unknown,
+III Weather‑KITTI and Weather‑NuScenes Datasets
+As introduced earlier, we propose our simulated adverse weather datasets, named Weather-KITTI and Weather-NuScenes, which are based on the SemanticKITTI[36] and nuScenes-lidarseg[37] datasets, respectively.
 To simulate three adverse weather conditions: snow, fog, and rain, we employ the LiDAR Snowfall Simulation (LSS) [33] for snow, the LiDAR Fog Simulation (LFS) [32] for fog, and the LiDAR Light Scattering Augmentation (LISA) [30] for rainy conditions.
-Table I presents the characteristics of the proposed dataset and existing publicly datasets.
-Compared to these existing adverse weather datasets, our proposed dataset provides a larger scale, covers a broader range of weather scenarios, offers complete original point cloud information, and has the capability to provide point-wise semantic and weather labels.
-III-A Adverse Weather Simulator
-LiDAR sensors emit laser pulses and measure travel distance based on the echoes received from targets, thereby capturing the precise shape of the targets[43].
-However, the LiDAR pulse system is susceptible to interference from scattering media such as particles of rain, fog, and snow, which can lead to shifts in measured distances, variations in echo intensity, and missing points, among other issues[44, 45, 46].
-The received signal power at a LiDAR’s receiver can typically be described using a linear system model[43], which is valid for non-elastic scattering.
-In this model, the range-dependent received signal power $P_{R}$ is modeled as the time-wise convolution between the time-dependent transmitted signal power $P_{T}$ and the time-dependent impulse response $H$ of the environment:
-[…]
-where $C_{A}$ is a system constant independent of time $t$, $R_{0}$ is the object distance and $c$ is the speed of light.
-1) Snow Model: In this work, we follow [40] to simulate these snow-corrupted data.
-This physically-based method samples snow particles in the 2D space and modify the measurement for each LiDAR beam in accordance with the induced geometry.
-For each point $p(x,y,z,i)$ from LiDARs, we calculate its snow-attenuated point $p_{\text{snow}}$ as follows:
-$R^{*}=\arg\max\left({P}_{R}\right)-c\frac{\tau_{H}}{2},\$ (2)
-$i_{\text{snow}}=P_{R}+i_{\max}f_{s}\left|f_{o}-\left(1-\frac{R^{*}}{R_{\max}}\right)\right|^{2},\$ (3)
-${{p_{\text{snow}}}}=\left(\frac{R^{*}}{R_{0}}\cdot(x,y,z),i_{\text{snow}}\right),$ (4)
-where $\tau_{H}$ is the LiDAR sensor’s half-power pulse width.
-$R^{*}$ and ${R_{\max}}$ represent the distances to the object at maximum received power $P_{R}$ and the furthest detectable object, respectively.
-$f_{s}$ is the focal slope and $f_{o}$ is the focal offset of the LiDAR sensor.
-As the snowfall rate $r_{s}$ increases, so does the number of particles in a point cloud frame, leading to denser snow particle accumulations in heavier snowfalls.
-Similar to [17], we define point clouds with changed spatial positions as the snow class.
-2) Fog Model: The physically-based fog simulation[41] is employed to transform clear LiDAR point clouds into their foggy counterparts.
-For each original point cloud, we calculate its fog-attenuated point $p_{\text{fog}}$ as follows:
-$i_{\text{hard }}={i}\cdot\exp({-2\alpha R_{0}}),$ (5)
-$i_{\text{soft}}={i}\cdot\frac{R_{0}^{2}}{\beta_{0}}\beta\cdot{i}_{tmp},$ (6)
-${{p_{\text{fog}}}}=\left\{\begin{array}[]{ll}\left(s\cdot n\cdot(x,y,z),{i_{\text{soft }}}\right),&\text{ if }{i_{\text{soft }}}>{i_{\text{hard }}}\\ \left(x,y,z,{i_{\text{hard }}}\right),&\text{ else.
-}\end{array}\right.$ (7)
-where $\alpha$ is the attenuation coefficient, $\beta$ denotes the fog backscattering coefficient, $\beta_{0}$ describes the differential reflectivity of the target.
-$i_{\text{soft }}$ and $i_{\text{hard }}$ represent the received responses for the soft target term and the hard target term, respectively.
-$s$ denotes scaling factor and $n$ is noise factor.
-${i}_{tmp}$ is the maximum response of the soft-target term.
-As $\beta$ increase, the fog rate rises, and the number of fog particles in the point cloud correspondingly increases.
-We similarly map point clouds with changes in three-dimensional spatial positions as the fog class.
-2) Rain Model: We employ the physics-based simulation model[42] to augment LiDAR scans for rain condition, which is based on a hybrid Monte-Carlo based approach.
-Specifically, the rain-distorted point cloud ${{p_{\text{rain}}}}$ is calculated as follows:
-[…]
-${i_{\text{rain}}}={i}\cdot\exp^{-2\alpha R_{\text{max }}}\cdot\min\left\{\left(\frac{D_{\text{max }}}{D_{b}\left(R_{\text{max }}\right)}\right)^{2},1\right\},$ (9)
-${{p_{\text{rain}}}}=\left(\gamma\cdot(x,y,z),i_{\text{rain}}\right),$ (10)
-where $\alpha$ is the extinction coefficient, $D$ is the particle diameter and $q_{e}$ is the extinction efficiency.
-$N$ denotes the particle density function.
-$D_{b}$ represents the beam diameter at the object distance $R_{\text{max }}$ and $\gamma$ is the rain-attenuated coefficient.
-Similarly, as the rain rate $r_{r}$ increases, the number of rain particles in a point cloud frame also correspondingly increases.
-We map point clouds with a rain-attenuated coefficient $\gamma$ not equal to 1 as the rain class.
+These simulation methods produce realistic weather effects, and they have demonstrated that perception models trained on their simulated data exhibit improved robustness under adverse weather conditions.
+Table I provides a detailed comparison between the proposed dataset and existing publicly available adverse‑weather datasets.
+Compared to these datasets, our dataset offers a larger scale, covers a broader range of weather scenarios, preserves complete point cloud information, and provides dense point‑wise semantic and weather annotations.
 ((a))
@@
 ((b))
-Figure 3: Semantic distribution of 3D points for all sequences in Weather-KITTI and Weather-NuScenes.
-Table II: Details of Weather Simulations. Each Weather
-Simulation Is Categorized Into Light, Moderate,
-and Heavy Levels Based on Different
-Weather Parameters.
-Weather Intensity Description
+((c))
+((d))
+Figure 3: Top: Semantic distribution of 3D points for all sequences in Weather-KITTI and Weather-NuScenes.
+Bottom: Distribution of weather-affected points across semantic classes under snow, fog, and rain.
+“Retained” denotes the percentage of points in that class unaffected by weather artifacts.
+Table II: Details of Weather Simulations.
+Weather Severity Description
 Snowfall Simulation Light snowfall rate $r_{s}=[0.5,1.0]$ mm/h
@@
 Heavy rain rate $r_{r}=[2.6,3.0]$ mm/h
-III-B Dataset Statistics
-Adverse weather primarily causes back-scattering and attenuation of LiDAR pulse transmissions, leading to significant shifts in both range and intensity for points in a LiDAR point cloud[43, 47].
-Figure 1 illustrates the semantic distributions of the same point cloud frame under three different weather conditions within our proposed Weather-KITTI dataset and their corresponding clean datasets.
-The weather noise points induced by the three weather conditions are primarily concentrated around the LiDAR sensor.
-Snowflakes contain ice crystals with varying densities, and in snowy conditions, each laser beam interacts with airborne particles, reflecting at different angles and considering potential occlusions.
-Compared to rain and fog, which consist of denser small water droplets, snowflakes result in more dispersed noise points.
-Table II delineates the severity levels of various weather-induced corruptions as configured in this study.
-In the snow simulation, the snowfall rate $r_{s}$ is set at [0.5, 1.0], [1.5, 2.0], and [2.5, 3.0] mm/h to represent light, moderate, and heavy snowfall, respectively.
-For fog simulations, the fog parameter $\beta$ is adjusted to [0.01, 0.05], [0.08, 0.14], and [0.18, 0.25] to reflect light, moderate, and heavy fog conditions.
-Similarly, in the rain simulation, the rain rate $r_{r}$ is set to [1.0, 1.5], [1.8, 2.4], and [2.6, 3.0] mm/h to simulate light, moderate, and heavy rainfall.
-As the severity of the weather increases, the extent of point cloud corruption proportionally intensifies.
-According to Table II, all LiDAR frames of Weather-KITTI and Weather-NuScenes are randomly synthesized with three different levels of weather severity in this study.
-This configuration allows us to investigate the generalizability of during models training.
-Simultaneously, we ensure that each level of weather severity has a similar number of LiDAR frames, guaranteeing the diversity of data across different weather scenarios.
-As depicted in Figure 2(a), the number of LiDAR frames for each weather category under the three weather severity levels in Weather-KITTI is approximately 14K frames.
-Similarly, as shown in Figure 2(b), the number of LiDAR frames for each weather category under each weather severity level in Weather-NuScenes is around 9K frames.
-Figure 4: Overview of the proposed denoising model, the TripleMixer network.
-The TripleMixer network comprises three main layers: the Geometry Mixer (GMX) Layer, the Frequency Mixer (FMX) Layer, and the Channel Mixer (CMX) Layer.
-These layers are detailed in Section IV.
-Weather-KITTI is a large-scale synthetic LiDAR dataset designed for adverse weather conditions, based on SemanticKITTI [38].
-It includes 22 sequences with a total of 130,656 LiDAR scans.
-Each point cloud is segmented into 21 semantic classes, with the semantic classes for snow, fog, and rain mapped to labels 110, 111, and 112, respectively, across all sequences.
-As SemanticKITTI provides semantic labels only for the first ten sequences, we also limit the availability of other semantic labels to the first ten sequences.
-Weather-NuScenes, derived from nuScenes-lidarseg [39], contains 84,390 LiDAR scans evenly divided into ten sequences.
-It includes 19 semantic classes and has similarly been enhanced with semantic classes for snow, fog, and rain across all sequences, mapped to labels 110, 111, and 112, respectively.
-Figure 3 illustrates the distribution of LiDAR points across semantic categories for all sequences in both Weather-KITTI and Weather-NuScenes.
-As depicted in the Figure 3(a), Weather-KITTI contains 139 million, 383 million, and 89 million LiDAR points for snow, fog, and rain noise, respectively.
-In comparison, as shown in Figure 3(b), Weather-NuScenes has 23 million, 112 million, and 6 million LiDAR points for snow, fog, and rain noise, respectively.
-This highlights the large scale and diverse weather conditions captured in our proposed datasets.
-IV Denoising Network
+III-A Dataset Statistics
+Figure 1 shows the semantic distributions of a single LiDAR frame captured under three different weather conditions from our Weather‑KITTI dataset, along with the corresponding clean reference frame.
+Weather‑induced noise points are predominantly concentrated near the LiDAR sensor, with snow producing more dispersed and irregular noise patterns than rain and fog.
+Table II presents the severity levels of various weather-induced corruptions configured in this study.
+Based on Table II, we simulated different severity levels (light, moderate, and heavy) of snow, fog, and rain.
+As weather severity increases, the degree of LiDAR point cloud corruption escalates accordingly.
+All LiDAR frames in the Weather‑KITTI and Weather‑NuScenes datasets were randomly synthesized across the three severity levels, with a comparable number of frames maintained for each level to ensure balanced data distribution.
+As illustrated in Figure 2(a), Weather-KITTI contains approximately 14K frames per weather type at each severity level, while Weather-NuScenes includes around 9K frames per category and severity level, as shown in Figure 2(b).
+In Weather‑KITTI, the proportions of degraded points at the light, moderate, and heavy severity levels across all weather classes (snow, fog, rain) are 17.55%, 30.85%, and 51.60%, respectively; In Weather‑NuScenes, the corresponding values are 14.51%, 37.21%, and 48.28%.
+Weather‑KITTI, based on SemanticKITTI [36], comprises 22 sequences with 130,656 LiDAR scans.
+Each point cloud is segmented into 21 semantic classes, with the weather-related classes for snow, fog, and rain mapped to labels 110, 111, and 112, respectively.
+As SemanticKITTI provides complete semantic labels only for the first ten sequences, we restrict the use of standard semantic labels accordingly.
+However, our dataset provides weather class annotations across all 22 sequences.
+Weather-NuScenes is derived from nuScenes-lidarseg [37] and includes 84,390 LiDAR scans evenly distributed across ten sequences.
+It covers 19 semantic classes, along with added labels for snow, fog, and rain using the same indices as in Weather-KITTI.
+The top row of Figure 3 shows the total number of LiDAR points per semantic class for all sequences in our datasets: Weather-KITTI contains 139 million, 383 million, and 89 million points corrupted by snow, fog, and rain noise, respectively.
+Similarly, Weather-NuScenes includes 23 million, 112 million, and 6 million corrupted points under the same conditions.
+The bottom row of Figure 3 presents the proportion of points affected by each weather artifacts per semantic class.
+Weather‑KITTI exhibits more pronounced weather effects on traffic‑related categories such as cars, trucks, and motorcycles, which are typically located near the LiDAR sensor and in central road regions where particles cause stronger occlusion, scattering, and reflection.
+In contrast, Weather‑NuScenes shows a more balanced distribution of weather corruption across classes, reflecting its diverse urban scenes and multiple LiDAR viewpoints.
+((a))
+((b))
+Figure 4: Spatial Distribution Characteristics of Noise Points in WADS and Our Dataset.
+Table III: Intensity Distribution Characteristics of Noise Points in WADS and Our Dataset.
+Intensity Range WADS (%) Ours (%)
+$\left[0.0,\ 0.3\right]$ 96.63 91.92
+$\left[0.3,\ 0.6\right]$ 3.01 7.54
+$\left[0.6,\ 1.0\right]$ 0.36 0.54
+III-B Dataset Comparison
+To assess the practical reliability of our simulated datasets, we compare them with WADS [18], which provides real‑world LiDAR sequences collected in snowy conditions with annotated noise labels.
+We randomly sample 1,000 scans from each dataset and analyze two key noise properties: spatial and intensity distributions.
+As illustrated in Figure 4, both datasets exhibit similar spatial characteristics.
+Approximately 90% of noise points are concentrated within a range of 3 to 10 meters from the LiDAR sensor, with a gradual decline beyond 10 meters and negligible presence past 20 meters.
+This consistency indicates that our simulations effectively reproduce real‑world occlusion and backscatter effects.
+Moreover, the intensity distribution of noise points, summarized in Table III, further demonstrates the similarity.
+Over 90% of noise points in both datasets fall within the low-intensity range [0.0, 0.3], indicating consistent degradation characteristics.
+Minor differences in the mid-range interval [0.3, 0.6] may be attributed to variations in LiDAR hardware or scene-specific factors, yet the overall distribution trend remains well aligned.
+These findings further validate that our synthetic weather corruption effectively replicates real-world noise patterns in both spatial and intesity properties.
+Figure 5: Overview of the proposed TripleMixer denoising network, which consists of three key layers: the Geometry Mixer (GMX), Frequency Mixer (FMX), and Channel Mixer (CMX).
+TripleMixer is trained solely using point-wise weather labels.
+IV TripleMixer Model
 IV-A Problem Statement
@@
 where $\theta$ denotes the model parameters to be optimized, aiming to minimize the difference between the prediction $\Phi({P_{i}};{\theta})$ and the ground truth labels $L_{i}$.
-IV-B The Overview of Network
+IV-B Overall Framework
 In this section, we propose the TripleMixer, a robust denoising model designed for removing 3D point cloud noises under adverse weather conditions.
-The architecture of our model is depicted in Figure 4.
-The TripleMixer consists of three primary layers: 1) the Geometry Mixer (GMX) Layer, which mixes the spatial geometric relationships of neighboring points, thereby enabling the network to discern crucial local structural features and maintain important geometric details; 2) the Frequency Mixer (FMX) Layer, employing multi-scale wavelet analysis to achieve a more comprehensive representation of both noise and normal points in the frequency domain, thereby enhances the model’s ability to discern fine-grained characteristics; and 3) the Channel Mixer (CMX) Layer, which enhances the overall feature representation by mixing contextual information across channels, improving the model’s accuracy and adaptability in complex environments.
+Its architecture is illustrated in Figure 5 and comprises three primary layers: 1) the Geometry Mixer (GMX) Layer, which mixes the spatial geometric relationships of neighboring points, thereby enabling the network to discern crucial local structural features and maintain important geometric details; 2) the Frequency Mixer (FMX) Layer, employing multi-scale wavelet analysis to achieve a more comprehensive representation of both noise and normal points in the frequency domain; and 3) the Channel Mixer (CMX) Layer, which enhances the overall feature representation by mixing contextual information across channels, improving the model’s accuracy and adaptability in complex environments.
 Specifically, in the GMX layer, we first voxelize the LiDAR point cloud in 3D space to facilitate downsampling, then select K neighboring points and employ the attentive pooling operation to effectively capture local geometric patterns.
@@
 The input features within the same voxel are meaned (downsampled), followed by the KNN neighborhood feature encoding, an attentive pooling operation, and a residual connection to obtain the output features.
+Figure 7: Architecture of the Frequency Mixer Layer.
+The Frequency Mixer Layer consists of three modules: the Triple Plane Projection (TPP) Block, the Lifting Wavelet (LWL) Block, and the Multi-resolution Feature Mixing (MFM) Block.
+It accepts input features with dimensions [H, V, C], where H and V represent height and width, and C represents the number of channels.
 IV-C Geometry Mixer (GMX) Layer
@@
 $F_{i}=Max\left(\left(MLP(f_{i})\right)\right)\oplus p_{i}.$ (5)
-Figure 6: Architecture of the Frequency Mixer Layer.
-The Frequency Mixer Layer consists of three modules: the Triple Plane Projection (TPP) Block, the Lifting Wavelet (LWL) Block, and the Multi-resolution Feature Mixing (MFM) Block.
-It accepts input features with dimensions [H, V, C], where H and V represent height and width, and C represents the number of channels.
 IV-D Frequency Mixer (FMX) Layer
@@
 Triple Plane Projection (TPP) Block.
-It’s aims to obtain the 2D features $x_{i}$ of the triple projection planes from the 3D features $F_{i}$ outputted by the GMX layer.
-To address the intensive computational demands of 3D convolution operations and the loss of height-related details in BEV projection, we introduce a novel approach by projecting the 3D point cloud along the X, Y, and Z axes onto three distinct 2D planes at varying resolutions, as illustrated in Figure 6.
+It aims to obtain the 2D features $x_{i}$ of the triple projection planes from the 3D features $F_{i}$ outputted by the GMX layer.
+To address reduce the computational complexity of 3D convolution operations and the loss of height-related details in BEV projection, we introduce a novel approach by projecting the 3D point cloud along the X, Y, and Z axes onto three distinct 2D planes at varying resolutions, as illustrated in Figure 7.
 The TPP block utilizes sparse tensor operations to efficiently handle data.
@@
 Consequently, the four sub-band components $\left[{x^{t}_{HL},x^{t}_{HH},x^{t}_{LL},x^{t}_{LH}}\right]\in\mathbb{R}^{H/2\times V/2\times C}$ at layer $t$ can be obtained as follows:
-[…]
-[…]
+$x^{t}_{HL},x^{t}_{HH}]=\mathit{VWD}\left(d^{t}\right),$ (8)
+$x^{t}_{LL},x^{t}_{LH}]=\mathit{VWD}\left(c^{t}\right),$ (9)
 where $L$ and $H$ denote low and high frequency information, respectively, while the first and second positions in each sub-band component’ indices correspond to the horizontal and vertical directions, respectively.
@@
 IV-F Loss Function
-To train our proposed model and ensure that the captured information is relevant to the point cloud denoising task, the overall loss function consists of three components: the cross-entropy loss, the lovász loss[48] and the wavelet regularization terms.
+To train our proposed model and ensure that the captured information is relevant to the point cloud denoising task, the overall loss function consists of three components: the cross-entropy loss, the lovasz loss[42] and the wavelet regularization terms.
 These components are structured as:
-$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\text{lovász }}+\mathcal{L}_{\mathrm{wr}}$ (23)
+$\mathcal{L}=\mathcal{L}_{\mathrm{ce}}+\mathcal{L}_{\text{lovasz}}+\mathcal{L}_{\mathrm{wr}}$ (13)
 $\mathcal{L}_{\mathrm{wr}}=\lambda_{1}\sum_{t=1}^{N}\left\|\mathcal{D}_{t}\right\|_{2}^{2}+\lambda_{2}\sum_{t=1}^{N}\left\|\mathcal{A}_{t}-\mathcal{A}_{t-1}\right\|_{2}^{2},$ (14)
-where $\mathcal{L}_{\mathrm{ce}}$ represents the cross-entropy loss, assessing the accuracy of noise predictions, and $\mathcal{L}_{\text{lovász}}$ denotes the lovász loss, which is designed for improving performance in tasks with imbalanced classes.
+where $\mathcal{L}_{\mathrm{ce}}$ represents the cross-entropy loss, assessing the accuracy of noise predictions, and $\mathcal{L}_{\text{lovasz}}$ denotes the lovasz loss, which is designed for improving performance in tasks with imbalanced classes.
 $\mathcal{L}_{\mathrm{wr}}$ represents the wavelet regularization terms, which include two term.
@@
 $\lambda_{1}$ and $\lambda_{2}$ are the wavelet regularization parameters used to adjust the strength of the regularization terms.
-$\mathcal{D}_{t}$ represents the mean of the detail sub-bands at decomposition level $t$, while $\mathcal{A}_{t}$ and $\mathcal{A}_{t-1}$ denote the mean of the approximation sub-bands at levels $t$ and level ${t-1}$ of the lifting wavelet, respectively.
-V Experiments
-In this section, we first introduce the experimental datasets and their division into training, validation, and testing subsets for model training.
-Next, we describe the implementation details and evaluation metrics of the experiments.
-Moreover, we conduct extensive performance comparison experiments to benchmark our model against state-of-the-art models.
-Finally, comprehensive ablation studies are performed to validate the contribution of each component of our model.
-V-A Datasets
-We conduct qualitative tests on three large-scale LiDAR datasets: WADS[17], as well as the Weather-KITTI and Weather-NuScenes datasets introduced in this paper.
-These datasets contain a variety of adverse weather conditions.
-The details of these datasets will be introduced as follows.
-WADS[17] contains LiDAR scans from 20 different sequences, with around 1300 LiDAR frames.
-The majority of these LiDAR scans are collected in snowy urban driving scenarios.
-Points in a LiDAR scan has been labeled into one of 22 semantic classes.
-The default labels for snow points are mapped as 110.
+$\mathcal{D}_{t}$ represents the mean of the detail sub-bands at decomposition level $t$, while $\mathcal{A}_{t}$ and $\mathcal{A}_{t-1}$ denote the mean of the approximation sub-bands at levels $t$ and level t1{t-1} of the lifting wavelet, respectively.
+V Denoising and Perception Benchmarks under Adverse Weather
+In this section, we establish a Denoising benchmark to evaluate the performance of our denoising model and introduce three downstream LiDAR perception benchmarks: Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD), to assess the generalization of state‑of‑the‑art perception models under adverse weather and the effectiveness of our denoising model as a preprocessing step.
+Notably, in all downstream benchmarks, our denoising model is trained in a supervised manner solely on our Weather‑KITTI and Weather‑NuScenes datasets using only point‑wise weather labels.
+Meanwhile, all perception models are directly tested on real‑world adverse‑weather datasets without any retraining or fine‑tuning.
+This design highlights the strong transferability of our datasets and the practical value of the proposed denoising models for real‑world applications.
+All experiments were conducted on an Ubuntu 18.04 with Intel Xeon Platinum 8280 CPUs, two Nvidia V100 GPUs, and 32 GB RAM.
+Table IV: Splitting of Training, Validation, and Testing Sets for Our Datasets.
+Weather-KITTI Weather-NuScenes
+Snow Scenarios Train: [01,02,08,17,19] [01,02,07,08]
+Val: [04,11,12,16] [04]
+Test:
+09,10,13,14,15,
+18,20,21] [00,03,05,06,09]
+Fog Scenarios Train:
+18,19,20] [01,02,03,07]
+Val: [06,07,09,11] [04]
+Test:
+12,13,14,15,21] [00,05,06,08,09]
+Rain Scenarios Train:
+12,16,17,21] [04,05,07,09]
+Val: [09,19,19] [08]
+Test:
+13,14,15,18,20] [00,01,02,03,06]
+V-A Denoising Benchmark
+Benchmark Setting: We conduct denoising experiments on the real‑world extreme‑weather dataset WADS [18] and our Weather‑KITTI and Weather‑NuScenes datasets.
+WADS[18] comprises 1,300 LiDAR scans across 20 sequences, collected in snowy urban driving scenarios.
 We divided the WADS dataset [^1] The training set is: [14,15,18,20,24,28,34,36,37].
@@
 into training, validation, and testing sets according to the protocol outlined in[14].
-To ensure a fair comparison for the point cloud denoising task under adverse weather conditions, we re-trained state-of-the-art models using our splits.
-Table III: Splitting of Training, Validation, and Testing Sets for Our Proposed Weather-KITTI and
-Weather-NuScenes Datasets.
-Weather-KITTI Weather-NuScenes
-Snow Scenarios Train: [00,02,08,17,19] [01,02,07,08]
-Val: [04,11,12,16] [04]
-Test: [01,03,05,06,07,
-09,10,13,14,15,
-18,20,21] [00,03,05,06,09]
-Fog Scenarios Train: [02,03,04,16,17,
-18,19,20] [01,02,03,07]
-Val: [06,07,09,11] [04]
-Test: [00,01,05,08,10,
-12,13,14,15,21] [00,05,06,08,09]
-Rain Scenarios Train: [01,02,03,04,06,
-12,16,17,21] [04,05,07,09]
-Val: [09,19,19] [08]
-Test: [00,05,07,08,10,
-13,14,15,18,20] [00,01,02,03,06]
-Weather-KITTI and Weather-NuScenes contain three different weather scenarios: snow, fog, and rain.
-Specifically, the Weather-KITTI dataset consists of 22 sequences for each weather scenario, while the Weather-NuScenes dataset includes 10 sequences per scenario.
-For model training, we divided the training, validation, and testing sets as detailed in Table V-A.
-V-B Implement Details
-In our experiments, we downsampled the point clouds by retaining only one point per 10 cm voxel.
-The resolutions for the 2D projections along the X, Y, and Z axes are set to [256, 256, 32], respectively.
-We optimized the model parameters using the AdamW optimizer[49] for 20 epochs with a weight decay of 0.005.
-The learning rate scheduler first performed a linear warmup over the initial 2 epochs, gradually increasing the learning rate from 0 to 0.001.
-After the warmup phase, it transitioned into a cosine annealing phase, where the learning rate decayed following a cosine curve, eventually reaching a minimum of $10^{-5}$ by the end of the last epoch.
-We set the batch size to 4 for each GPU and employed various data augmentation techniques to enhance model robustness, including random rotations, scaling, and random flipping along the XY plane.
-Additionally, we used stochastic depth with a layer drop probability of 0.2.
-The loss is the sum of the cross-entropy loss, the lovász loss and the wavelet regularization terms.
-The regularization parameters $\lambda_{1}$ and $\lambda_{2}$ are set to 0.1 for all the experiments.
-All results are produced on an Ubuntu 18.04 system equipped with two Intel® Xeon® Platinum 8280 CPUs, two Nvidia V100 GPUs, and 32 GB of RAM.
-V-C Evaluation Metrics
-Following prior work[22, 21, 20], we evaluate the point cloud denoising task using $precision=\frac{TP}{{TP+FP}}$, $recall=\frac{TP}{{TP+FN}}$, F1 score, and mean Intersection over Union (mIoU) as quantitative metrics.
-Specifically, the F1 score and mIoU are defined as follows:
+To ensure a fair comparison, we re-trained state-of-the-art models using our splits.
+For Weather‑KITTI and Weather‑NuScenes, the dataset partitions for training, validation, and testing are provided in Table V.
+During training, point clouds were downsampled to a 10 cm voxel grid and projected onto the X, Y, and Z planes at [256, 256, 32] resolutions.
+Models were trained for 30 epochs using AdamW [43] with a weight decay of 0.005, a batch size of 4 per GPU, and stochastic depth (drop probability 0.2).
+The learning rate warmed up linearly to 0.001 over 2 epochs, then decayed to $10^{-5}$ with cosine annealing.
+The loss combined cross‑entropy, Lovasz, and wavelet regularization with $\lambda_{1}=\lambda_{2}=0.1$.
+Following prior work [14, 21, 20], we evaluate the point cloud denoising task using precision, recall, F1 score, and mean Intersection‑over‑Union (mIoU) as quantitative metrics, where $precision=\frac{TP}{TP+FP}$ and $recall=\frac{TP}{TP+FN}$.
+Specifically, the F1 score and mIoU are defined as
 ${F_{1}}=2\times$ (15)
 ${mIoU}=\frac{TP}{{TP+FP+FN}}$ (16)
-Where $TP$ refers to the correctly identified noise points, $FP$ refers to the non-noise points incorrectly identified as noise, and $FN$ refers to the noise points that were not correctly identified.
-V-D Quantitative Result
-Results On WADS: We compare our model with the current state-of-the-art models, including both statistical and deep learning approaches.
-Specifically, SOR[16], ROR[16], DSOR[17], and DROR[30] are statistical methods.
-SalsaNext[32] and Cylinder3D[33] are general point cloud segmentation methods that we have adapted for binary classification to enable comparison.
-WeatherNet[20], 4DenoiseNet[21], and 3D-OutDet[22] are specialized point cloud denoising networks.
+where $TP$ denotes correctly identified noise points, $FP$ denotes non‑noise points incorrectly identified as noise, and $FN$ denotes noise points that were not correctly identified.
+Results On WADS: We compare our model with the current state‑of‑the‑art models, with a primary focus on point cloud denoising networks.
+Specifically, WeatherNet [20], 4DenoiseNet [21], and 3D‑OutDet [14] are specialized point cloud denoising networks.
+In addition, SOR [17], ROR [17], DSOR [18], and DROR [25] are statistical methods, while SalsaNext [44], Cylinder3D [45], and LSK3DNet [46] are point cloud segmentation methods that we have adapted for binary classification to enable comparison.
 Table LABEL:wads presents a comprehensive comparison of denoising performance on the WADS[18] dataset, evaluating precision, recall, F1 score, and mIoU.
-The WADS[17] dataset represents real-world conditions, making it a robust benchmark for performance assessment.
-As illustrated in Table LABEL:wads, while Cylinder3D[33] achieves the highest precision with a score of 97.12, our model closely follows with a minimal difference of approximately 0.06.
-More importantly, our model outperforms all existing methods in other critical metrics, achieving the highest recall, F1 score, and mIoU with scores of 93.93, 94.97, and 90.73, respectively.
+As illustrated in Table LABEL:wads, while LSK3DNet[46] achieves the highest precision with a score of 97.16, our model surpasses all existing methods in other critical metrics, achieving the highest recall, F1 score, and mIoU with scores of 93.93, 95.13, and 90.73, respectively.
 These results underscore the superior efficacy of our model in point cloud denoising, particularly in complex real-world applications.
-Table IV: Denoising Performance Results on WADS Datasets. The Proposed TripleMixer Achieves the Best mIoU.
+Table V: Denoising Performance Results on WADS Datasets.
 Method Precision↑ Recall↑ F1↑ mIoU↑
+SalsaNext[44] 74.16 93.50 82.71 70.52
+Cylinder3D[45] 97.12 92.61 94.81 90.13
+LSK3DNet[46] 97.16 92.70 94.88 90.25
 ROR[17] 17.13 91.80 29.15 17.06
@@
 DROR[25] 69.84 90.10 78.68 64.05
-SalsaNext[32] 74.16 93.50 82.71 70.52
-Cylinder3D[33] 97.12 92.61 94.81 90.13
 WeatherNet[20] 96.69 81.24 88.28 79.02
@@
 TripleMixer (Ours) 96.38 93.93 95.13 90.73
-Table V: Denoising Performance Results On The Proposed Weather-KITTI Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
+Figure 8: Qualitative results of the Denoising benchmark on the WADS [18], Weather‑KITTI (WK), and Weather‑NuScenes (WN) datasets.
+Semantic point colors follow the scheme in Figure 1, with red points indicating weather‑induced noise.
+Table VI: Denoising Performance On The Proposed Weather-KITTI Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios.
 Method Average Snow Scenarios Fog Scenarios Rain Scenarios
 mIoU Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑
-DROR[30] 56.24 44.18 97.36 60.77 43.66 66.05 92.74 78.63 64.79 61.47 96.68 74.11 60.28
 SalsaNext[44] 89.58 97.40 92.83 95.06 90.58 97.19 94.59 95.87 92.07 97.75 87.82 92.52 86.08
 Cylinder3D[45] 94.28 99.13 96.39 97.74 95.58 99.78 92.92 96.23 92.73 96.79 97.58 97.18 94.52
+LSK3DNet[46] 94.66 98.95 96.85 97.87 95.94 99.60 93.75 96.51 93.16 97.20 97.62 97.40 94.88
+DROR[25] 56.24 44.18 97.36 60.77 43.66 66.05 92.74 78.63 64.79 61.47 96.68 74.11 60.28
 WeatherNet[20] 79.15 83.28 97.20 89.70 81.33 89.16 86.64 87.88 78.38 78.49 98.81 87.48 77.75
-4DenoiseNet[21] 82.85 85.23 97.08 90.77 83.10 89.90 93.95 91.88 84.98 83.02 96.34 89.18 80.48
-3D-OutDet[22] 90.55 98.14 95.15 96.63 93.48 99.61 85.26 91.88 84.97 94.12 98.98 96.48 93.21
+4DenoiseNet[21] 81.85 85.23 97.08 90.77 83.10 89.90 93.95 91.88 81.98 83.02 96.34 89.18 80.48
+3D-OutDet[14] 91.55 98.14 95.15 96.63 93.48 99.61 85.26 91.88 87.97 94.12 98.98 96.48 93.21
 TripleMixer (Ours) 96.31 99.23 97.18 98.16 96.46 99.03 96.50 97.74 95.61 97.66 99.15 98.40 96.87
-Table VI: Denoising Performance Results On The Proposed Weather-NuScenes Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
+Table VII: Denoising Performance Results On The Proposed Weather-NuScenes Datasets. Average mIoU Represents the Mean mIOU Across Three Weather Scenarios: Snow, Fog, and Rain.
 Method Average Snow Scenarios Fog Scenarios Rain Scenarios
 mIoU Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑ Precision↑ Recall↑ F1↑ mIoU↑
-DROR[30] 42.93 38.36 92.74 54.27 37.24 40.36 98.68 57.28 40.14 51.71 98.89 67.91 51.41
 SalsaNext[44] 73.12 80.48 75.23 77.76 63.21 93.79 84.21 88.74 79.76 82.74 90.89 86.62 76.40
 Cylinder3D[45] 95.41 98.85 95.56 97.18 94.51 98.94 96.37 97.64 95.38 99.24 97.06 98.14 96.34
+LSK3DNet[46] 96.04 98.55 97.20 97.86 94.96 99.10 97.70 98.39 96.08 99.35 98.00 98.66 97.10
+DROR[25] 42.93 38.36 92.74 54.27 37.24 40.36 98.68 57.28 40.14 51.71 98.89 67.91 51.41
 WeatherNet[20] 71.73 72.96 85.40 78.69 64.87 84.97 80.54 82.69 70.49 80.44 99.06 88.78 79.83
@@
 TripleMixer (Ours) 97.21 97.32 98.37 97.84 95.78 99.37 98.02 98.69 97.42 99.72 98.71 99.21 98.43
-Figure 7: Visualization of Denoising Performance Results on Weather-KITTI and Weather-NuScenes Datasets. Our TripleMixer demonstrates consistently stable performance across different scenarios.
-Figure 8: Qualitative results among the proposed TripleMixer and state-of-the-art denoising models[21, 22, 33] on the WADS[17], the proposed Weather-KITTI and Weather-NuScenes datasets.
-To highlight the differences, the correct / incorrect predictions are painted in gray / red, respectively.
-Best viewed in colors and zoomed-in for details.
-Results On Weather-KITTI: In the previous section, we demonstrated the superior performance of our denoising model on the real-world WADS[17] dataset.
-To further validate the good generalizability of our model, we conducted additional denoising comparison experiments on our proposed Weather-KITTI and Weather-NuScenes datasets.
+Results On Weather-KITTI: To further validate the good generalizability of our model, we conducted additional denoising comparison experiments on our proposed Weather-KITTI and Weather-NuScenes datasets.
 For a thorough evaluation, we selected the most recent and best-performing models from Table LABEL:wads for further comparative analysis.
-As shown in Table LABEL:Weather-KITTI, we compared our model with six state-of-the-art methods across three scenarios in the Weather-KITTI dataset: Snow Scenarios, Fog Scenarios and Rain Scenarios.
-Our model achieved the highest average mIoU of 96.31 across these three scenarios, surpassing popular general semantic segmentation methods such as Cylinder3D[33], as well as the latest denoising models like 4DenoiseNet[21] and 3D-OutDet[22].
-Specifically, our model’s average mIoU on the Weather-KITTI dataset is 2.2% higher than Cylinder3D[33], 16.2% higher than 4DenoiseNet[21], and 5.9% higher than 3D-OutDet[22].
-It is noteworthy that our model consistently achieved the highest mIoUs of 96.46, 95.61, and 96.87 across Snow, Fog, and Rain scenarios, respectively, while DROR[30], Cylinder3D[33], and SalsaNext[32] achieved the highest Recall or Precision in the Snow, Fog, and Rain scenarios, respectively.
-This demonstrates the robustness of our model in point cloud denoising on the Weather-KITTI dataset.
-Results On Weather-NuScenes: We also compared the denoising performance of our model with six state-of-the-art methods across three scenarios in the Weather-NuScenes dataset.
+As shown in Table LABEL:Weather-KITTI, we compared our model with seven state-of-the-art methods across three scenarios in the Weather-KITTI dataset: Snow Scenarios, Fog Scenarios and Rain Scenarios.
+Our model achieved the highest average mIoU of 96.31, outperforming general semantic segmentation methods such as LSK3DNet [46] and advanced denoising models like 4DenoiseNet [21] and 3D‑OutDet [14].
+Specifically, our model’s average mIoU exceeded LSK3DNet, 4DenoiseNet, and 3D‑OutDet by 1.7%, 16.2%, and 5.9%, respectively.
+Notably, our model achieved the highest mIoUs of 96.46, 95.61, and 96.87 in Snow, Fog, and Rain scenarios, respectively, whereas Cylinder3D [45], and SalsaNext [44] attained the highest Precision in individual scenarios.
+This likely reflects their conservative filtering or task‑specific feature modeling, which improves precision by suppressing false positives but often reduces recall.
+In contrast, our Frequency Mixer (FMX) module performs multi‑scale frequency decomposition through a learnable lifting‑wavelet design, enabling the selective retention of task‑relevant high‑frequency details.
+This design prioritizes structural completeness over aggressive filtering, thereby striking a better balance between precision and recall and ensuring robust and consistent denoising across diverse weather conditions.
+Results On Weather-NuScenes: We also compared the denoising performance of our model with seven state-of-the-art methods across three scenarios in the Weather-NuScenes dataset.
 As shown in Table LABEL:Weather-NuScenes, our TripleMixer model achieved the highest average mIoU of 97.21 across these three scenarios.
@@
 These results demonstrate that our model exhibits superior performance in the Weather-NuScenes dataset across three different adverse weather scenarios compared to other denoising models.
-We further present the mIoU performance of our model and three other denoising models across six scenarios in the Weather-KITTI and Weather-NuScenes datasets, as shown in Figure 7.
-Our model not only achieved the highest mIoU but also exhibited the smallest variance in mIoU accuracy across the six scenarios, with a variance of only 0.96.
-This indicates that our denoising model has strong generalization capability and is the most robust compared to the other denoising models, effectively denoising across diverse weather conditions.
-Visualization: We visualized our model’s prediction results for point cloud denoising performance on WADS[17], Weather-KITTI, and Weather-NuScenes, compared with those of the state-of-the-art models, including: 4DenoiseNet[21], 3D-OutDet[22], and Cylinder3D[33].
-As illustrated in Figure 8, it is evident that our TripleMixer model delivers more accurate predictions, achieving optimal denoising results across different weather scenarios compared to the other three denoising models.
-VI Discussions
-VI-A Ablation Study:
-In this section, we perform thorough ablation studies on the point cloud denoising task to investigate the effect of each components, projection plane resolution, wavelet decomposition levels and loss functions of our TripleMixer model.
-All ablation experiments are conducted on the real-world WADS[17] dataset, with same experimental settings as in Section V.
-Table VII: Ablation Results of Each Component. GMX: Geometry Mixer Layer, FMX: Frequency Mixer Layer, and CMX: Channel Mixer Layer.
+Figure 8 visualizes denoising results on WADS [18], Weather‑KITTI, and Weather‑NuScenes, showing that TripleMixer consistently produces more accurate and cleaner reconstructions than 3D‑OutDet [14] and LSK3DNet [46] across various weather conditions.
+Table VIII: Benchmarking Results for the SS Task (%).
+All SS models are pretrained on SemanticKITTI [36] and inferred on SemanticSTF [23].
+Denoising models are trained on Weather-KITTI and used as preprocessing during inference.
+Methods mIoU car bi.cle mt.cle truck oth-v. pers. bi.list mt.list road parki. sidew. oth-g. build. fence veget. trunk terra. pole. traf.
+$S$ Preprocessing
+SphereFormer [47] Baseline 11.26 19.12 0.00 0.00 13.83 4.12 0.99 6.75 0.00 22.28 4.69 15.80 4.43 25.53 13.27 24.23 4.38 37.74 14.71 1.99
+4DenoiseNet [21] 11.03 18.24 0.00 0.00 13.42 3.92 1.00 2.17 0.00 22.01 4.72 16.25 4.74 25.33 14.68 23.97 4.50 38.15 14.57 1.98
+3D-OutDet [14] 11.87 23.26 0.00 0.00 15.94 3.58 1.64 1.54 0.00 24.44 4.68 17.41 5.73 25.04 13.99 27.25 5.30 38.57 15.20 2.04
+TripleMixer (Ours) 13.71 27.19 0.00 0.00 21.20 4.07 0.21 5.68 0.00 26.72 4.76 19.68 8.85 23.97 33.64 22.86 4.31 40.95 14.48 1.84
+SFPNet [48] Baseline 11.14 14.36 0.00 0.00 11.60 0.00 0.23 5.29 0.00 38.06 4.33 13.71 5.09 17.19 12.06 30.01 4.97 37.98 14.86 1.92
+4DenoiseNet [21] 11.07 14.01 0.00 0.00 10.89 0.00 0.24 4.55 0.00 38.18 4.12 13.73 5.08 17.26 12.39 29.73 4.96 38.07 15.10 1.95
+3D-OutDet [14] 12.12 23.51 0.00 0.00 11.27 0.00 0.34 4.14 0.00 43.87 6.45 14.83 5.81 17.30 11.76 30.38 4.88 38.37 15.42 2.01
+TripleMixer (Ours) 13.23 21.33 0.06 0.00 15.95 0.00 0.36 4.91 0.00 42.35 8.37 15.74 6.58 19.88 18.61 37.01 5.05 38.00 15.39 1.87
+PTv3 [49] Baseline 14.58 15.27 1.13 0.00 3.56 0.40 0.21 0.00 37.25 44.10 12.05 9.95 6.46 25.30 25.70 37.46 8.63 36.97 10.10 2.46
+4DenoiseNet [21] 14.53 14.88 1.12 0.00 3.60 0.48 0.21 0.00 37.25 44.07 11.86 9.92 6.43 25.42 25.95 37.57 8.70 37.00 9.22 2.48
+3D-OutDet [14] 16.40 19.87 2.14 0.00 9.58 2.48 0.21 0.00 40.56 44.51 11.30 9.87 8.48 25.28 30.03 40.60 8.54 42.18 12.12 2.79
+TripleMixer (Ours) 17.18 23.64 2.80 0.00 19.88 8.26 1.74 0.00 42.87 44.06 11.36 9.91 9.36 29.09 28.53 36.93 9.26 36.93 9.87 1.98
+Figure 9: Qualitative results of the SS benchmark under adverse weather.
+The “Denoising” row presents outputs of different denoising models, while the “Semantic Segmentation” row shows PTv3[49] results on the corresponding denoised point clouds.
+Red points denote weather‑induced noise, and blue points represent the “car” class.
+V-B Semantic Segmentation (SS) Benchmark
+Benchmark Setting: We evaluate the downstream SS models on the real‑world adverse‑weather SemanticSTF [23] dataset, which provides semantic annotations for 19 categories consistent with SemanticKITTI.
+To comprehensively assess performance, we select three recent state‑of‑the‑art semantic segmentation models: SphereFormer [47], SFPNet [48], and Point Transformer v3 (PTv3) [49].
+In our experiments, they pretrained on SemanticKITTI under clear‑weather conditions are directly evaluated on SemanticSTF.
+Furthermore, to examine the impact of denoising on SS models performance, we incorporate three specialized denoising networks: 4DenoiseNet[21], 3D‑OutDet [14], and our proposed TripleMixer, all trained in a supervised manner on the our Weather‑KITTI datasets and directly generalized to SemanticSTF.
+Evaluation follows standard protocols, computing the Intersection‑over‑Union (IoU) for each class and reporting the mean IoU (mIoU) across all classes.
+Benchmark Results: Table LABEL:LSS_BENCH presents the benchmarking results for the SS task under real-world adverse weather on the validation set of SemanticSTF [23] dataset.
+The baseline denotes direct application of SS models without denoising preprocessing.
+Across all segmentation backbones, including SphereFormer [47], SFPNet [48], and PTv3 [49], the proposed TripleMixer consistently yields the highest mIoU, demonstrating significant improvements over both the baseline and other denoising methods.
+Specifically, when integrated with SphereFormer, TripleMixer achieves an mIoU of 13.71, improving by 2.45 points over the baseline.
+Similar gains are observed with SFPNet and PTv3, where TripleMixer improves the mIoU by 2.09 and 2.60 points, respectively.
+On average, TripleMixer achieves a 2.38-point improvement across all three backbones, significantly surpassing the performance of 3D-OutDet, which only provides an average gain of 1.13 points.
+Notably, 4DenoiseNet, which demonstrates limited denoising effectiveness on our proposed datasets, fails to improve real-world segmentation performance and even leads to performance degradation compared to the baseline.
+At the category level, TripleMixer delivers substantial gains in key semantic classes, improving IoU by 7.81 for car and 9.35 for truck over the baseline, averaged across the three segmentation models.
+It also enhances performance in challenging structural categories such as buildings and fences.
+These results underscore the robustness and practical utility of TripleMixer and our dataset for improving LiDAR semantic segmentation in real‑world conditions.
+We visualize the SS benchmark under adverse weather in Figure 9 to illustrate denoising results in real‑world scenarios and the subsequent semantic segmentation performance, focusing on the key semantic class “car”.
+Compared with 3D‑OutDet, TripleMixer achieves more effective denoising and enhances segmentation performance under real‑world extreme conditions.
+Table IX: Benchmarking Results for the PR Task.
+All PR models are trained on the KITTI [35] and inferred on Boreas (BR) [41] and Weather-KITTI (WK).
+Denoising models, trained on Weather-KITTI, serve as preprocessing during inference.
+Method mRS WK Snow WK Fog WK Rain BR Easy BR Hard
+PR Preprocessing R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑ R@1↑ R@1%↑ F1↑
+OT [50] Baseline 0.15 0.03 0.26 0.07 0.09 0.33 0.17 0.17 0.52 0.29 0.01 0.11 0.02 0.00 0.09 0.02
+4DenoiseNet [21] 0.22 0.03 0.24 0.06 0.21 0.58 0.37 0.25 0.61 0.40 0.01 0.09 0.02 0.00 0.06 0.01
+3D-OutDet [14] 0.27 0.07 0.38 0.13 0.22 0.69 0.42 0.27 0.62 0.43 0.02 0.13 0.03 0.01 0.11 0.03
+TripleMixer (Ours) 0.31 0.09 0.43 0.17 0.31 0.78 0.47 0.29 0.65 0.44 0.04 0.15 0.06 0.03 0.14 0.05
+CVTNet [51] Baseline 0.28 0.03 0.34 0.07 0.29 0.50 0.46 0.19 0.61 0.33 0.14 0.21 0.25 0.09 0.14 0.15
+4DenoiseNet [21] 0.33 0.02 0.30 0.03 0.31 0.48 0.46 0.43 0.83 0.60 0.14 0.20 0.25 0.10 0.15 0.17
+3D-OutDet [14] 0.37 0.04 0.42 0.09 0.31 0.50 0.47 0.49 0.86 0.65 0.17 0.23 0.26 0.12 0.17 0.19
+TripleMixer (Ours) 0.39 0.06 0.44 0.11 0.33 0.54 0.50 0.51 0.89 0.67 0.16 0.29 0.28 0.13 0.24 0.21
+LPSNet [52] Baseline 0.30 0.02 0.18 0.04 0.09 0.77 0.16 0.29 0.74 0.46 0.13 0.28 0.26 0.10 0.26 0.19
+4DenoiseNet [21] 0.29 0.02 0.22 0.05 0.03 0.77 0.06 0.28 0.72 0.44 0.16 0.28 0.27 0.10 0.24 0.19
+3D-OutDet [14] 0.33 0.02 0.27 0.06 0.10 0.85 0.17 0.32 0.75 0.49 0.18 0.29 0.28 0.11 0.27 0.20
+TripleMixer (Ours) 0.37 0.03 0.29 0.06 0.14 0.83 0.24 0.37 0.79 0.54 0.15 0.36 0.29 0.11 0.34 0.23
+V-C Place Recognition (PR) Benchmark
+Benchmark Setting: We selected distinct sequences from the real‑world Boreas [41] dataset and our Weather‑KITTI dataset.
+In Boreas, the sequence recorded on 2021-11-02-11-16 is used as the database, while the sequence from 2021-01-26-11-22, which features heavy snowfall, serves as the query.
+Boreas Easy employs only the forward traversal, whereas Boreas Hard includes both directions to introduce greater viewpoint variation.
+For Weather‑KITTI, we use the snow 00, fog 07, and rain 00 as query sets, with their corresponding clear-weather sequences serving as the database.
+The benchmark includes three leading place recognition models: OverlapTransformer (OT)[50], CVTNet[51], and LPSNet [52], all pretrained on the KITTI dataset and directly evaluated on Boreas and Weather‑KITTI.
+The denoising models and training configurations are the same as those used in the SS benchmark.
+For the PR evaluation, a match is considered a true recognition if the pose distance is less than 5 meters for the Weather-KITTI, and less than 15 meters for the Boreas dataset.
+We follow standard metrics widely adopted in prior work [51], including Recall@1 (R@1), Recall@1% (R@1%), and the F1 score.
+Additionally, we introduce the Mean Robust Score (mRS) to quantify model robustness under adverse weather conditions, defined as:
+${\text{mRS}}=\frac{1}{N}\sum\frac{\text{M}_{\text{w}}}{\text{M}_{\text{c}}},\quad\text{M}\in\{\text{R@1},\text{R@1\%},\text{F1}\},$ (17)
+where $\text{M}_{\text{w}}$ and $\text{M}_{\text{c}}$ denote the PR evaluation metrics under adverse and clean weather conditions, respectively.
+$N$ represents the total number of evaluated metrics across all test sequences.
+A higher value of mRS reflects greater robustness of the PR model when transitioning from clean to adverse weather conditions
+Benchmark Results: Table LABEL:LPR_BENCH presents the benchmarking results for the PR task under adverse weather conditions on the Boreas and Weather-KITTI datasets.
+Across all test sequences, TripleMixer consistently achieves the highest mRS values, demonstrating its effectiveness in enhancing the robustness of PR models under adverse weather conditions.
+For OT, TripleMixer improves the mRS from the baseline value of 0.15 to 0.31.
+Similarly, for CVTNet and LPSNet, TripleMixer achieves mRS values of 0.39 and 0.37, respectively, exceeding both the baseline and other denoising models.
+TripleMixer achieves an average improvement of 56% across three PR models under adverse weather conditions.
+Notably, it consistently enhances place recognition performance even in the challenging Boreas Hard scenarios.
+These results highlight the effectiveness of both TripleMixer and the proposed Weather-KITTI dataset in improving the robustness of downstream PR models against real-world adverse weather conditions.
+Figure 10 visualizes the result of PR benchmark.
+Most noise points are concentrated near the LiDAR center.
+Our TripleMixer achieves effective denoising on both the real‑world Boreas dataset and our Weather‑KITTI dataset, thereby improving PR accuracy.
+Figure 10: Qualitative results of the PR benchmark under adverse weather.
+The PR results show queries processed by different denoising methods and their top‑1 retrieved matches using CVTNet [51].
+Brown boxes indicate incorrect retrieval, while blue boxes denote correct retrievals.
+Table X: Benchmarking Results for the OD Task (%).
+All OD models are trained on the KITTI dataset [35] and inferred on CADC [29].
+Denoising models, trained on Weather-NuScenes, are employed as preprocessing during inference.
+Methods mAP Car (3D) Car (BEV) Pedestrian (3D) Pedestrian (BEV)
+OD Preprocessing Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑ Easy↑ Med↑ Hard↑
+TED-S [53] Baseline 12.18 21.85 20.07 15.76 24.16 21.13 18.65 4.26 4.17 3.47 4.54 4.46 3.72
+4DenoiseNet [21] 11.91 22.31 20.03 14.88 25.20 21.77 17.51 4.24 3.12 2.81 4.35 3.71 3.01
+3D-OutDet [14] 12.63 22.33 20.05 16.66 25.18 22.74 18.50 5.02 4.12 3.64 5.23 4.31 3.80
+TripleMixer (Ours) 13.58 24.35 22.06 18.65 27.17 23.73 19.49 5.24 4.20 3.81 5.54 4.73 4.02
+PG-RCNN [54] Baseline 12.53 22.36 19.11 16.67 24.91 22.51 19.89 5.36 5.05 1.99 5.41 5.09 2.04
+4DenoiseNet [21] 12.25 22.60 19.13 16.72 25.81 22.30 19.73 4.31 4.17 1.59 4.33 4.21 1.84
+3D-OutDet [14] 13.43 23.79 19.93 16.11 26.84 23.09 18.96 6.28 6.16 3.80 6.29 6.18 3.84
+TripleMixer (Ours) 14.14 24.90 20.23 17.81 28.32 23.43 20.82 6.42 6.39 4.23 6.45 6.43 4.27
+VoxT-GNN [55] Baseline 10.02 15.34 14.24 11.93 19.91 18.78 15.38 4.37 4.16 3.71 4.46 4.25 3.79
+4DenoiseNet [21] 9.70 15.81 13.84 11.58 20.40 18.23 14.78 3.63 3.48 3.25 4.08 3.70 3.56
+3D-OutDet [14] 10.53 16.59 14.94 12.55 21.18 19.31 15.81 4.62 4.18 4.01 4.85 4.26 4.09
+TripleMixer (Ours) 12.33 19.41 16.37 14.13 23.99 20.93 18.09 6.41 6.38 4.96 6.14 6.12 4.93
+V-D Object Detection (OD) Benchmark
+Benchmark Setting: For the OD benchmark, we use the real‑world CADC [29] dataset, which provides object‑level annotations for 3D object detection in a KITTI‑style format, ensuring compatibility with standard evaluation protocols.
+We evaluate three advanced 3D object detection models, TED‑S [53], PG‑RCNN [54], and VoxT‑GNN [55], using their pretrained versions on the KITTI dataset and directly test them on CADC dataset.
+Consistent with the SS benchmark, we employ the same denoising models, trained on Weather‑NuScenes and evaluated on CADC, leveraging their shared 32‑beam LiDAR configuration.
+For evaluation, we follow the KITTI 3D object detection protocol, reporting the mean Average Precision (mAP) over 40 recall positions (R40).
+We adopt IoU thresholds of 0.7 for cars and 0.5 for pedestrians, while excluding cyclists due to their absence in CADC.
+Results are reported across three difficulty levels: easy, moderate, and hard.
+Benchmark Results: The benchmarking results for the OD task under real-world adverse weather conditions on the validation set of CADC [29] dataset, as summarized in Table LABEL:LOD_BENCH.
+Evaluations are conducted for both 3D and Bird’s Eye View detection across car and pedestrian categories.
+The results demonstrate that incorporating TripleMixer as a preprocessing module consistently improves both 3D and BEV detection performance under real-world adverse weather conditions across all OD baselines, including TED-S [53], PG-RCNN [54], and VoxT-GNN [55], with an average performance gain of 16%.
+Specifically, when applied to PG-RCNN, it increases the overall mAP from 12.53 to 14.14 and improves 3D detection for the Car category under hard conditions from 16.67 to 17.81.
+Similarly, with VoxT-GNN, it boosts the mAP from 10.02 to 12.33 and enhances Pedestrian 3D detection under hard conditions from 3.71 to 4.96.
+In contrast, 4DenoiseNet fails to improve detection performance, similar to its results in the SS task.
+Its limited denoising ability hinders performance on detail-sensitive tasks.
+These results show that TripleMixer benefits from our Weather-NuScenes dataset, enabling more effective spatial restoration and consistently improving OD performance under real-world extreme weather across different backbones and categories.
+Figure 11 illustrates the performance of OD models with different denoising preprocessing under adverse weather.
+Models without denoising exhibit numerous false positives and missed detections.
+Compared with 3D‑OutDet, our TripleMixer produces cleaner point clouds, enabling OD models to achieve higher detection accuracy and reducing both false positives and missed detections.
+Figure 11: Qualitative results of the OD benchmark under adverse weather.
+The “Denoising” rows display outputs from different denoising models, and the subsequent rows present object detection results on the corresponding denoised point clouds.
+Red boxes indicate ground‑truth objects, green boxes denote predicted objects.
+Turquoise circles highlight incorrect detections, and pink circles indicate missed detections.
+Table XI: Ablation Results for Each Layer of TripleMixer.
 Experiment GMX FMX CMX Metric(%)
@@
 IV 96.38 93.93 95.13 90.73
-Table VIII: Ablation Results of Different XY Directions Projection Resolution When Z Direction Projection Resolution is set to 32
+Table XII: Ablation Results for XY Projection Resolutions with Z Resolution Fixed at 32.
 XY-Resolution Precision↑ Recall↑ F1↑ mIoU↑
@@
 [512,512] 96.25 98.85 95.03 90.53
-Table IX: Ablation Results of Different Z Direction Projection Resolution When XY Directions Projection Resolution is set to [256,256]
+Table XIII: Ablation Results for Z Projection Resolutions with XY Resolutions Fixed at [256, 256]
 Z-Resolution Precision↑ Recall↑ F1↑ mIoU↑
@@
 [48] 96.46 93.59 95.00 90.48
-Table X: Ablation Results of Different Wavelet Decomposition Levels in the Frequency Mixer Layer
+Table XIV: Ablation Results for Wavelet Decomposition Levels in FMX Layer
 Methods Precision↑ Recall↑ F1↑ mIoU↑
@@
 w/ Wavelet-3 96.16 93.80 94.96 90.41
-[…]
+Table XV: Ablation Results of Loss Functions.
 […]
@@
 IV 95.13 90.73
-Table XII: Semantic Segmentation Performance Comparison on WADS Datasets.
-The Row Marked with $\Delta$ Represents the Performance Difference of Our Denoising Model TripleMixer for Downstream Semantic Segmentation Tasks.
-Methods mIoU
-car
-truck
-other-vehicle
-person
-road
-parking
-sidewalk
-other-ground
-building
-fence
-other-structure
-vegetation
-trunk
-terrain
-pole
-traffic-sign
-other-objec
-acc-snow
-SalsaNext[32] 14.32 45.98 4.84 0.00 0.00 50.56 4.21 6.53 2.47 46.85 2.40 1.05 47.35 0.34 0.31 6.12 7.46 0.29 29.91
-SalsaNext[32] + TripleMixer 15.61 50.22 1.20 0.45 0.93 53.83 14.67 8.79 7.92 54.97 3.86 0.50 43.25 0.29 0.35 7.08 11.00 1.41 20.22
-$\Delta$ +1.29 +4.24 -3.64 +0.45 +0.93 +3.27 +10.46 +2.26 +5.45 +8.12 +1.46 -0.55 -4.10 -0.05 +0.04 +0.96 +3.54 +1.12 -9.69
-Cylinder3D[33] 19.40 61.88 0.00 0.00 0.00 47.45 0.15 4.46 0.17 72.01 16.03 1.01 51.76 0.00 0.00 24.40 14.42 0.36 54.16
-Cylinder3D[33] + TripleMixer 21.03 66.03 0.00 0.26 0.48 46.90 2.20 5.99 0.57 79.13 14.47 1.37 57.80 0.18 0.24 31.19 17.92 1.62 52.28
-$\Delta$ +1.63 +4.15 0.00 +0.26 +0.48 -0.55 +1.05 +1.53 +0.40 +7.12 -1.56 +0.36 +6.04 +0.18 +0.24 +6.79 +3.50 +1.26 -1.88
-PolarNet[50] 19.28 59.10 0.03 0.00 0.00 48.25 0.01 0.14 0.20 72.30 3.84 0.01 61.94 0.00 0.00 19.94 21.77 0.39 59.27
-PolarNet[50] + TripleMixer 20.36 60.21 0.08 0.12 0.06 59.38 0.01 0.01 0.04 68.40 13.29 0.42 59.08 0.22 0.26 27.45 21.78 2.32 53.23
-$\Delta$ +1.08 +1.11 +0.05 +0.12 +0.06 +11.13 0.00 -0.13 -0.16 -3.90 +9.45 +0.41 -2.86 +0.22 +0.26 +7.51 +0.01 +1.93 -6.04
-Effects of Each Component.
-To testify the effectiveness of each component of TripleMixer, we conduct an extensive ablation study and list the result in Table VII.
-As shown in Table VII, Experiments I-III represent the accuracy of our TripleMixer model after sequentially removing the GMX Layer, the FMX Layer, and the CMX Layer, respectively.
+VI Discussion
+VI-A Ablation Study
+This section presents ablation studies to evaluate the impact of key components in the TripleMixer model, including three primary layers, projection plane resolution, wavelet decomposition levels and loss functions.
+All ablation experiments are conducted on the real-world WADS[18] dataset, with same experimental settings as in Section V.
+Effects of Each Layer.
+To testify the effectiveness of each Layer of TripleMixer, we conduct an extensive ablation study and list the result in Table XI.
+Experiments I-III represent the accuracy of our TripleMixer model after sequentially removing the GMX Layer, the FMX Layer, and the CMX Layer, respectively.
 From Experiments I and IV, we observe that the GMX Layer contributes significantly, resulting in F1 and mIoU scores of 0.49 and 0.90, respectively.
@@
 Effects of Projection Plane Resolution.
-In the Frequency Mixer Layer, the 3D geometric features are projected onto three 2D planes along the X, Y, and Z axes.
-An adaptive wavelet decomposition is then performed on these projected 2D planes to extract multi-scale features.
-The initial projection resolution of each 2D plane influences the effectiveness of the multi-scale feature extraction.
-In this part, we conduct ablation experiments to assess the impact of different projection plane resolutions on the performance of our TripleMixer model.
-Due to the different horizontal and vertical emission ranges of LiDAR, we set different projection resolutions for the XY directions and the Z direction.
-We first set the projection resolution for the Z direction to 32 in order to select the optimal projection resolution for the XY directions.
-Table LABEL:plane_reso_xy shows the impact of different projection resolutions in the XY directions on the model’s denoising performance, with the Z direction projection resolution set to 32.
-As shown in Table LABEL:plane_reso_xy, the optimal mIoU accuracy is achieved when the projection resolution in the XY directions is set to [256,256].
-Both excessively high resolutions such as [512,512] and excessively low resolutions such as [64,64] result in suboptimal performance.
-Therefore, we selected the resolution of [256,256] for the XY directions.
-We further verified the impact of different resolutions in the Z direction on model accuracy.
-The results are shown in Table LABEL:plane_reso_z, where we experimented with five different resolutions for the Z direction.
-As indicated in Table LABEL:plane_reso_z, with the XY directions’ resolution fixed, the optimal mIoU accuracy is achieved when the Z direction resolution is set to 32.
-Similar to the results for the XY directions, excessively high or low resolutions in the Z direction also resulted in poorer accuracy.
-Therefore, in our practical experiments, we chose the projection resolution of [256,256,32] for the XYZ directions.
+In the FMX layer, 3D geometric features are projected onto three 2D planes along the X, Y, and Z axes.
+Adaptive wavelet decomposition is then applied to extract multi-scale features from each plane.
+The initial resolution of these projections plays a key role in determining the quality of the extracted features.
+To evaluate its impact, we conduct ablation studies on different projection resolutions.
+Considering the asymmetric LiDAR scanning ranges, we separately investigate the XY and Z directions.
+First, we fix the Z-direction resolution at 32 and explore various resolutions in the XY plane.
+As shown in Table LABEL:plane_reso_xy, the highest mIoU is achieved at [256, 256], while both lower ([64, 64]) and higher ([512, 512]) resolutions lead to performance drops.
+Based on this, we fix the XY resolution to [256, 256] and examine different Z-direction resolutions.
+Table LABEL:plane_reso_z shows that a resolution of 32 again yields the best performance, with other values resulting in suboptimal accuracy.
+These findings confirm that the projection resolution of [256, 256, 32] provides the best balance between feature granularity and computational efficiency, and is adopted in all practical experiments.
 Effects of Wavelet Decomposition Levels.
-In this part, we investigate the impact of varying levels of wavelet decomposition in the Frequency Mixer Layer on our model’s performance.
-With the projection resolution set to [256,256,32] for the XYZ directions, we select a maximum wavelet decomposition level of 3.
-This choice is due to the fact that if the wavelet decomposition level exceeds 3, the resolution in the Z direction becomes $\leq 2$, which hinders the extraction of effective information.
-Consequently, we analyze the model’s performance across a range of wavelet decomposition levels, from none to the maximum level of 3.
-As shown in Table LABEL:wavelet, not employing any wavelet decomposition levels results in the poorest model accuracy, highlighting the critical role that wavelet transformations play in optimizing model performance.
-Additionally, utilizing the maximum number of decomposition levels also leads to suboptimal performance, as excessive wavelet decomposition fails to extract more effective information.
-Our TripleMixer achieves the best overall accuracy when the decomposition level is set to 2.
+We evaluate the impact of different wavelet decomposition levels in the Frequency Mixer (FMX) on model performance.
+With the projection resolution fixed at [256, 256, 32] along the XYZ axes, we set the maximum wavelet decomposition level to 3, as higher levels reduce the Z-direction resolution to less than or equal to 2, limiting the extraction of meaningful features.
+We examine decomposition levels ranging from 0 to 3.
+As shown in Table LABEL:wavelet, omitting wavelet decomposition yields the lowest accuracy, underscoring its importance in feature extraction.
+While moderate decomposition improves performance, excessive decomposition introduces information loss.
+The best results are obtained when the wavelet decomposition level is set to 2, confirming its effectiveness in balancing detail preservation and multi-scale representation.
 Effects of Loss Functions.
-Our model’s loss function primarily consists of three components: the cross-entropy loss, the Lovász loss, and the wavelet regularization terms.
+Our model’s loss function primarily consists of three components: the cross-entropy loss, the lovasz loss, and the wavelet regularization terms.
 We use the cross-entropy loss as the baseline to test the impact of different loss functions on model accuracy.
-As shown in Table XI, after adding only the Lovász loss, the mIoU and F1 scores of our TripleMixer model improved by 0.37 and 0.21, respectively.
+As shown in Table XV, after adding only the lovasz loss, the mIoU and F1 scores of our TripleMixer model improved by 0.37 and 0.21, respectively.
 Additionally, after incorporating only the wavelet regularization terms, the model’s mIoU and F1 scores increased by 0.26 and 0.14, respectively.
 When using the combination of all three loss functions, our denoising model achieved the highest mIoU and F1 scores.
-VI-B Experiments on Segmentation Tasks
-In this section, we apply the proposed denoising model to LiDAR-based downstream perception tasks.
-We conducted point cloud semantic segmentation experiments on the WADS dataset, both with and without our denoising model.
-These experimental results demonstrate the generalization ability of our proposed method and its effectiveness on LiDAR point cloud processing.
-The performance of the point cloud semantic segmentation task for SalsaNext[32], Cylinder3D[33] and PolarNet[50] models before and after removing snow with our TripleMixer model is shown in Table LABEL:wads_denoise, where $\Delta$ indicates the performance differences resulting from the snow removal.
-As illustrated in Table LABEL:wads_denoise, with the assistance of our model, the semantic segmentation performance of the SalsaNext[32], Cylinder3D[33], and PolarNet[50] models has significantly improved.
-The performance enhancements are quantified as increases of 1.29%, 1.63%, and 1.08%, respectively.
-Notably, before the snow removal by TripleMixer, none of the three models were capable of detecting the ’other-vehicle’ and ’person’ categories.
-However, with the integration of our model, there is a marked improvement in the detection accuracy for these categories.
-This demonstrates that our model is beneficial and versatile for LiDAR-based downstream perception tasks.
-VII Conclusion and Future Work
-In this work, we first introduce two large-scale adverse weather point cloud datasets, Weather-KITTI and Weather-NuScenes, featuring a wide range of weather scenarios, complete original LiDAR information, and point-wise weather and semantic labels.
-These datasets provide robust data support for further research into adverse weather conditions.
-Subsequently, we propose a novel point cloud denoising model, the TripleMixer network, which integrates spatial geometric features, frequency domain features, and multi-channel contextual information to achieve high denoising accuracy.
-This model demonstrates optimal performance on both real-world datasets and our newly proposed datasets.
-Finally, we validate the significant performance enhancement of our denoising model for downstream point cloud tasks, specifically semantic segmentation, under adverse weather conditions.
-In future work, we aim to incorporate additional sensors, such as cameras and IMUs, to further explore the perception capabilities of multi-sensor fusion in extreme weather conditions.
+Table XVI: Comparison of Model Complexity, Runtime, and Peak Memory Usage.
+Method Parameters (M) Runtime (ms) Memory (GB)
+Cylinder3D [45] 49.52 127 0.92
+LSK3DNet[46] 28.84 89 0.78
+DSOR [18] 253
+DROR [25] 199
+4DenoiseNet[21] 0.57 96 2.47
+3D-OutDet[14] 0.01 82 0.33
+TripleMixer (Ours) 1.02 61 0.21
+VI-B Efficiency Comparisons
+Table XVI compares the model complexity, runtime, and peak memory usage during inference across different methods.
+General point cloud segmentation models adapted for denoising, such as Cylinder3D [45] and LSK3DNet [46], have substantially higher parameter counts of 49.52M and 28.84M and memory footprints of 0.92GB and 0.78GB.
+Traditional statistical denoising methods DSOR [18] and DROR [25] are computationally expensive with long runtimes.
+3D‑OutDet[14] has the smallest parameter count of 0.01M but a relatively high runtime due to additional k‑NN preprocessing.
+4DenoiseNet [21] shows the highest peak memory usage of 2.47GB as it aggregates spatiotemporal features from multiple frames.
+In contrast, our TripleMixer achieves a favorable balance with a runtime of 61ms and the lowest peak memory usage of 0.21GB, demonstrating high efficiency.
+VII Conclusion
+In this work, we proposed TripleMixer, a robust plug‑and‑play denoising network that can be seamlessly integrated into LiDAR perception pipelines.
+We also introduced two large‑scale simulated LiDAR datasets, Weather‑KITTI and Weather‑NuScenes, encompassing diverse adverse weather conditions.
+Building on these contributions, we established four benchmarks: Denoising, Semantic Segmentation (SS), Place Recognition (PR), and Object Detection (OD), providing a unified framework for evaluating denoising models and their impact on downstream perception tasks.
+Our model was trained exclusively on the proposed datasets, and all perception models were directly evaluated on real‑world adverse‑weather datasets without retraining.
+Extensive experiments demonstrated that TripleMixer achieves state‑of‑the‑art denoising performance and substantially improves SS, PR, and OD accuracy under real‑world adverse weather, underscoring the strong generalization capability of both the proposed model and datasets.
+Future Work.
+While the multi‑branch mixer design of TripleMixer effectively suppresses noise, its relatively high parameter count may hinder deployment in resource‑constrained settings.
+Future work will focus on developing a more lightweight denoising network.
+Furthermore, while synthetic weather data provides valuable scalability and controllability, it may not fully capture the subtle complexities of real‑world weather conditions.
+We will explore more realistic simulation techniques and expand real‑world adverse‑weather datasets to further improve model generalization.